### v1

# Using $k$-medoids for distributed approximate similarity search with arbitrary distances

Abstract
This paper presents GMASK, a general algorithm for distributed approximate similarity search that accepts any arbitrary distance function. GMASK requires a clustering algorithm that induces Voronoi regions in a dataset and returns a representative element for each region. Then, it creates a multilevel indexing structure suitable for large datasets with high dimensionality and sparsity, usually stored in distributed systems. Many similarity search algorithms rely on $k$-means, typically associated with the Euclidean distance, which is inappropriate for specific problems. Instead, in this work we implement GMASK using $k$-medoids to make it compatible with any distance and a wider range of problems. Experimental results verify the applicability of this method with real datasets, improving the performance of alternative algorithms for approximate similarity search. In addition, results confirm existing intuitions regarding the advantages of using certain instances of the Minkowski distance in high-dimensional datasets.
Keywords Approximate similarity search  $\cdot$ k-medoids  $\cdot$ arbitrary distance function
1 Introduction
The goal of similarity search is to find elements from a dataset, represented by feature vectors in a multidimensional space, similar to a given query object according to a particular similarity or dissimilarity measure. Numerous data indexing methods, collectively known as access methods, have been proposed to speed up this task. Nevertheless, the results remain strongly conditioned by the problems inherent to the representation of datasets in high-dimensionality and high-sparsity spaces. These algorithms are frequently conceived to work with data having spatial components or represented in a vector space with few dimensions. However, they become unusable in high-dimensional spaces, where many features characterise each element due to the Curse of Dimensionality [1]. As a result, the distance between elements tends to become increasingly uniform, making it challenging to differentiate between them based solely on their distances or when the number of dimensions significantly exceeds the quantity of data points. Moreover, several studies indicate that in datasets with over 20 dimensions, exact similarity search methods rarely outperform exhaustive scanning.
Owing to this constraint, novel approaches have emerged in recent years to swiftly identify enough related items within a reasonable time frame compared to linear scanning, often referred to as approximate similarity search. One of the newest approaches is MASK (Multilevel Approximate Similarity search with k-means) [2], a novel indexing method that involves an unconventional application of the $k$-means partitioning algorithm [3, 4] to create a multilevel index structure for approximate similarity search.
Traditionally, the input to $k$-means is always a fixed number of $k$ prototype points or centroids. Then, following an iterative process, the dataset is divided into $k$ groups (Voronoi cells), assigning each point to its closest centroid according to a given distance function, so that each centroid represents all of the elements of that cluster. In contrast, leveraging current powerful computing capabilities, $k$-means can be forced to generate a high number of centroids that inundate the dataset. Then, these centroids become a new dataset to be grouped at the upper level, repeating this process until the complete construction of the search tree is achieved. This bottom-up creation of the multilevel index enables the distribution of this task among different computing nodes, which also store distinct data partitions.
Although it employs an innovative approach that capitalises on the benefits offered by contemporary computing systems and distributed platforms, the performance of this indexing method remains significantly impacted by the inherent constraints of the $k$-means algorithm. The choice of the distance function significantly impacts the outcome of $k$-means clustering, as it determines how the similarity between data points is calculated and can lead to different outcomes in terms of the shape and separation of the clusters.
Until now, experiments conducted with $k$-means have relied on the Euclidean distance to calculate relative distances between points, following the conventional inclination to associate the concept of distance with that used in geospatial contexts. Nonetheless, relying on the Euclidean distance as the default option in indexing structures and algorithms when working with high-dimensional data, despite its widespread use, can lead to sub-optimal results. This is because, in high-dimensional spaces, the discriminating power of the $L_{p}$ norm decreases for higher values of $p$ due to their specific features (such as dispersion, outliers, and concentration of points around certain values), leading to poor differentiation between proximate and distant elements.
This paper introduces GMASK (General Multilevel Approximate Similarity Search with $k$-medoids), a novel algorithm that facilitates the incorporation of arbitrary distances by employing data partitioning algorithms that induce Voronoi regions in a dataset and yield a representative element, exemplified by $k$-medoids. The simple introduction of $k$-medoids turns GMASK into a versatile algorithm capable of adapting to the characteristics of any dataset (high dispersion, high dimensionality, outliers, etc.) and seek for the optimal results. This approach allows using multiple $\ell_{p}$ norms or even other distance functions, making it compatible with a broader spectrum of problems. To comprehensively evaluate its performance, an extensive series of experiments will be conducted across datasets varying in type, size, dimensionality, and sparsity.
This paper adheres to the following structure. Section “Background and related work” reviews previous research related to similarity search. Section “Methodology” describes GMASK, emphasising the potential for enhancing the algorithm’s performance by exploring the application of other distance functions beyond the Euclidean one, like other Minkowski distances. The findings of four empirical experiments carried out over real-world datasets with quite diverse attributes are presented in Section “Experimental results”. These tests compare GMASK performance with other well-known algorithms, such as FLANN and PyNNdescent. Finally, Section “Conclusions and future work” recaps the main conclusions of this work and suggests promising lines for future work.
1.1 Contributions
The potential for achieving improved results, coupled with the lack of prior empirical studies that validate this proposal and the need for learning paradigms applicable to a wider range of dissimilarity measures [5], justifies our efforts to explore alternative Minkowski distances for data indexing. This endeavour is aimed at challenging the misconception we have identified and fostering a deeper understanding of the significant impact of using alternative distance metrics when working with high-dimensional data.
  • GMASK allows the use of different clustering algorithms for the bottom-top construction of the index tree, such as $k$-medoids, a broader form of the $k$-means algorithm that offers parameterisation with several distance metrics, in particular Minkowski distances.
  • Evaluating the performance of different $L_{p}$ norms or Minkowski distances for approximate $k$-nearest neighbour search. This assessment involves testing GMASK against other state-of-the-art methods to illustrate that the selection of a suitable distance function, while considering the features of the target dataset, has a significant impact on performance when searching for reasonably similar regions.
2 Background and related work
This section provides minimal background on the concepts and strategies underlying similarity search methods, including the description of current similarity search algorithms.
2.1 Similarity search
Similarity search, often referred to as proximity search [6, 7, 8], is a cornerstone for applications in many different fields, such as databases [9], information retrieval [10], computer vision [11] and bioinformatics [12, 13], among others. It aims to find elements that are similar (close) to a given query object, according to some measure of similarity or, conversely, dissimilarity.
Thus, given a query object $q\in\mathbb{X}$, the similarity search tackles some of the following scenarios, represented in Figure 1 [7]:
  • Point query: Finding elements within $\mathbb{X}$ that have exactly the same feature values as $q$.
  • Range query: Identifying a subset of elements $\{o_{i}\}\in\mathbb{X}$ whose feature values do not exceed a given similarity threshold $r$ around $q$ such that $\forall\{o_{i}\},\delta(o_{i},q)\leq r$.
  • Nearest-neighbour query: Finding elements in $\mathbb{X}$ whose features are the closest to the feature vector representing $q$. In this case, the goal might be to retrieve the single element (set of $k$ elements) closest to $q$, known as the nearest neighbour ($k$-nearest neighbours) and denoted by $NN_{q}$ ($k-NN_{q}$). Identifying the $k$-nearest neighbours of a given object is one of the most widely used classification methods in machine learning.
Figure 1: Conceptual representation of several types of query searches.
Many techniques for data indexing, collectively known as access methods, can speed up these queries. These methods create an index structure that partitions the domain $\mathbb{X}$ into different regions based on a distance function. Then, we can leverage these regions to solve similarity queries. There is a wide range of access methods, including exact approaches, approximate solutions and variants tailored for distributed computing systems.
2.2 Algorithms for similarity search
Exact similarity search methods identify the completely accurate result set for a query by using the absolute positions of objects to group elements and search for similar instances. These methods are categorised as Point Access Methods (PAM) for point-based elements and Spatial Access Methods (SAM) for elements with spatial dimensions, such as lines, polygons or higher-dimensional polyhedra. The general acronym SAM is commonly used in the research literature to refer to both categories of indexing strategies in vector spaces [10]. These approaches are most effective with elements represented in a relatively small number of dimensions, making Curse of Dimensionality one of the biggest challenges in data science today.
Metric Access Methods (MAM) provide a general framework for similarity search problems within metric spaces [6, 8], exploiting the triangle inequality to partition the metric space into subspaces, effectively filtering out portions of the dataset that cannot contain valid results. On the one hand, MAM exact search can be expensive in terms of computational time and updating the index structure due to dynamic data. On the other hand, MAM indexes have serious problems in efficiently reducing the search for candidate elements in a query when dealing with high-dimensional data. There are many MAM for queries in metric spaces, some of which use clustering algorithms, such as $k$-means [4] or $k$-medoids [14], to recursively partition the entire set of elements creating a hierarchical structure (tree) of nested clusters.
Approximate indexing methods are solutions that strike a balance between accuracy and speed, making them suitable for both vector and metric spaces [15]. Although these methods may introduce some errors into the results, they also provide a faster response time by calculating fewer distances. As a result, these methods prove feasible in scenarios where exact methods cannot provide a fast response, such as in high-dimensional problems or large datasets, where exact methods default to a full scan. Some of the most widely used exact algorithms are compared and summarised in [16], which also introduces a comprehensive taxonomy to classify them according to relevant characteristics.
Rather than computing the exact result set for a query, approximate similarity search algorithms adopt various strategies (e.g., modifying the representation space, reducing the number of pair distances to compute) to identify a reduced set of candidate items that are likely to be the most similar. FastMap, introduced in [17], is one of the first algorithms to be developed for approximate similarity search in metric spaces. It is based on a method to obtain a fast projection of elements in the original dataset into a user-defined $k$-dimensional space, preserving the distances between pairs of elements in order to achieve an approximate response through dimensionality reduction. In this category of algorithms, MASK [2] emerged as a novel indexing approach characterised by a multilevel design tailored for approximate similarity search. It features an innovative utilisation of the $k$-means partitioning algorithm to quickly reduce the number of checked regions when searching for results.
2.3 Using different distance functions
The dissimilarity value captures the notion of closeness between two objects, which is interpreted as a distance in a suitable space. A distance function that avoids introducing any correlation between features, such as the Minkowski distance [18] or $\ell_{p}$ norm, can quantify this value [17]. The Manhattan or city block distance ($\ell_{1}$ norm), the Euclidean distance ($\ell_{2}$ norm), and the Chebyshev distance ($\ell_{\infty}$ norm) are typical instances of this family. As many multidimensional indexing structures were initially proposed in the context of spatial systems, many indexing structures and algorithms have commonly relied on the Euclidean distance, derived from its conventional use in two- or three-dimensional spatial applications.
However, several negative consequences and challenges arise when dealing with high dimensional problems, an effect widely known as the Curse of Dimensionality. As the number of dimensions increases, the volume of the space grows exponentially, and the density of points in the space decreases. This sparsity poses a challenge in accurately measuring distances and similarities between data points, since proximity becomes less meaningful. Consequently, points can be very far apart, even though they are close according to a metric such as Euclidean distance. This effect increases the computational complexity of algorithms and operations performed in high-dimensional spaces, often resulting in slower and less efficient computations. For this reason, several authors have pointed out that, due to the Curse of Dimensionality, it is very costly to find the exact nearest neighbour in high-dimensional Euclidean spaces [19, 20].
As a result, although Euclidean spaces reflect our natural conception of space, this natural assumption is challenging to extrapolate when dealing with high-dimensional data. Moreover, traditional clustering algorithms, such as $k$-means, often specify the Euclidean distance as the only appropriate metric without considering other possibilities. This choice brings in a self-imposed limitation that may produce results far from the best [21, 22, 5].
In recent years, several surveys [23, 24, 25] have addressed the evaluation of various metrics for similarity searches across spaces of diverse topology, a burgeoning topic offering insights into the most suitable similarity measures for different problems and contributing to the ongoing development of new methods for similarity search.
Therefore, we aim to introduce GMASK, an algorithm that, being based on a generalisation of earlier concepts previously introduced by MASK, signifies a substantial departure from the latter. GMASK enables the integration of arbitrary distances through the use of data partitioning algorithms that generate Voronoi regions and assign a prototype for each of these regions. Specifically, the utilisation of the $k$-medoids algorithm allows for the incorporation of multiple $\ell_{p}$ norms, facilitating parametrisation with different distance metrics and the execution of validation experiments. Our findings demonstrate the feasibility of employing diverse distances beyond the Euclidean metric, highlighting the potential for achieving superior results in data indexing, contingent upon the specific characteristics of the dataset.
3 Methodology
This section describes the methodology developed for data indexation using arbitrary distance functions. First, it illustrates the GMASK algorithm as a general approximate indexing method accepting any distance function. Hereafter, it delves into the characteristics of Minkowski distances, emphasising the inherent limitations in $k$-means that justify the implementation of GMASK. Consequently, we justify adopting $k$-medoids, an alternative clustering method that leverages the opportunities provided by Minkowski distances to overcome previous restrictions derived from using the Euclidean distance. The design of experimental tests is described in Subsections “Performance evaluation”, “Benchmark algorithms for similarity search”, and “Datasets”, including metrics to assess performance, state-of-the-art algorithms for comparison, and the datasets selected to gauge effectiveness.
3.1 GMASK
As mentioned above, there is a close connection between machine learning and similarity search indexing. Firstly, the foundations for designing data access methods rely on familiar concepts in machine learning classification, such as dissimilarity spaces and distance functions. Secondly, clustering methods like $k$-means or $k$-medoids are used in many of these algorithms to create indexing structures for similarity search and improve search performance. Besides, they are also affected by the same limitations, such as the Curse of Dimensionality and the challenge of identifying the actual intrinsic dimensionality governing some problems to make them solvable.
The typical choice in partitioning clustering methods is to find a low value of $k$ that generates different groups in the dataset. GMASK follows an unconventional approach for clustering, generating a high number of $k$ representative points for each level. The only requirement in GMASK is using a clustering method that generates Voronoi regions and returns a representative point for each of them. In this implementation, GMASK uses the $k$-medoids clustering method, inducing a Voronoi tessellation and returning an element representing each region. Thus, generating a large number of representative points, GMASK summarises the underlying data points with finer detail. In the same way, this indexing method applies to any dissimilarity space, including the particular case of metric spaces or vector spaces. Thanks to the $k$-medoids adaptation in GMASK, this algorithm presents an important improvement over MASK: it is able to work with different distance functions. This modification brings GMASK the capability of adaptation to the specific characteristics of any dataset. In the “Experimental results” section, we show that this unusual application of clustering has attractive properties and yields superior performance, even with high-dimensional datasets residing in feature spaces that are typically sparse.
At this point, it is essential to note that the goal of GMASK is not to cluster data but to use the partitioning clustering method to construct a multilevel index to find elements. With more powerful computational infrastructures and large memory and storage capacity becoming prevalent, the advantage of this method lies in that previous restrictions on the number of clusters at each level become less relevant. In this new scenario, we can exploit the effect of introducing a high number of prototype points into the dataset, leveraging the capabilities of the computing infrastructure.
Figure 2 illustrates the design principles of GMASK. At the lowest level, we have the data points in $\mathcal{X}$ to be indexed. The blue boxes in the diagram represent different data partitions or groups that could be stored on various computing nodes. The novelty lies in creating the multilevel structure of $k$ medoids, following a bottom-up approach. The first level of medoids (identified as $n_{1}$ in Figure 2) summarises the data points at the lowest level so that points assigned to each cluster are represented by its prototype (medoid). The second level (denoted $n_{2}$) corresponds to medoids representing the prototypes computed at the first level, $n_{1}$. The same strategy is repeated, creating additional prototype layers that summarise the immediately lower group of medoids. The procedure stops when the number of medoids at the top level of this index structure becomes manageable.
Figure 2: Conceptual representation of the updated multilevel structure approach created by GMASK.
The process of constructing this multilevel index structure depends on two initial parameters:
  • The group length quantifies the number of elements in each data partition at the lowest level, configured according to the available computing power (and the total number of nodes in distributed computing systems). For the baseline, we assume that all partitions are of the same size, although this condition is not strictly necessary.
  • The number of medoids calculated for each group. At this point, the question of selecting the most optimal number of medoids has not yet been addressed, but it is not very important. Contrary to the standard procedure, our approach suggests that, for information indexing, one should use as many medoids as the computational infrastructure can afford.
The clustering at the first level reduces the initial number of data points by a particular proportion, the data summarisation ratio, which is determined by the ratio between the group length and the number of medoids. This ratio also determines the number of layers of the multilevel index and the stop condition when the number of medoids at the top layer becomes lower than the group length.
Each medoid is a prototypical point representing all the elements assigned to its cluster. The number of points assigned to each cluster may vary depending on their distance from the corresponding medoid. For this reason, regions with higher data density will also tend to receive more medoids. By generating a high number of medoids, we can capture finer details of the underlying data, allowing us to index not only high-density regions but also outliers and clusters of elements distinct from the main concentrations of points. Algorithm 1 describes the construction of this multilevel index structure.
Algorithm 1 Multilevel index construction in GMASK
1:function Clustering($distFunc,nPrototypes,points$)
2:     $\triangleright$ A clustering method returning Voronoi regions
3:     $\triangleright$ and a prototype point representing each region
4:end function
5:
6:procedure GMASK($data,lengthGroup,nPrototypes$)
7:     $ngroups\leftarrow lengthData/lengthGroup$;
8:     $vector\leftarrow$ split($data,nGroups$); $\triangleright$ Divide dataset into nGroups
9:     $\triangleright$ Initialise lists $layerPoints$, $layerLabels$, and variable $idLayer$
10:     while $ngroups\geq 1$ do
11:         $\triangleright$ Initialise groupsPoints, groupsLabels and points
12:         for $idGroup=1,nGroups$ do
13:              $points\leftarrow vector[idGroup]$;
14:              $groupsPoints.add($clustering($distFunc,nPrototypes,points$));
15:              $groupsLabels.add(clusters.labels)$;
16:         end for
17:         $layerPoints.add(groupsPoints)$;
18:         $layerLabels.add(groupsLabels)$;
19:         $ngroups\leftarrow length(groupsPoints)/lengthGroup$;
20:         if $ngroups\geq 1$ then
21:              $vector\leftarrow$ split($layerPoints,nGroups$);
22:         end if
23:         $idLayer\mathrel{+}=1$;
24:     end while
25:     $layers\leftarrow idLayer-1$;
26:end procedure
Once built according to the rules described above, the index is ready to accept point queries, which systematically search for all points in the original dataset, seeking a perfect match. To solve any query using this multilevel index, we start at the top level of the medoids hierarchy, steering the search according to the distance between the medoids at each level and the target query object $q\in\mathcal{X}$.
3.2 Dissimilarity measures
A dissimilarity space [26] defines the rules for measuring the distance between elements represented in a feature space. As discussed in Section “Similarity search”, similarity captures the notion of closeness between two objects. Euclidean vector spaces are widely applied in previous research, as they reflect and intuitive interpretation of a space: vectors are elements that can be added, multiplied by a scalar or projected onto each other.
Being $\mathbb{X}$ a set of elements, the function
$\delta:\mathbb{X}\times\mathbb{X}\to\mathbb{R}$ (1)
returns a real number $\delta(a,b)$ that shall be used to quantify the difference between two or more elements contained on this set, denoting a measure of distance.
Distance functions have the following properties:
  • $(p1)\quad\forall a,b\in\mathbb{X},\delta(a,b)\geq 0$ positiveness,
  • $(p2)\quad\forall a,b\in\mathbb{X},\delta(a,b)=\delta(b,a)$ symmetry,
  • $(p3)\quad\forall a\in\mathbb{X},\delta(a,a)=0$ reflexivity,
  • and in most cases
  • $(p4)\quad\forall a,b\in\mathbb{X},a\neq b\to\delta(a,b)>0$ strict positiveness.
If it also satisfies
  • $(p5)\quad\forall a,b,c\in\mathbb{X},\delta(a,b)\leq\delta(a,c)+\delta(c,b)$, commonly known as triangle inequality,
$\delta$ is indeed a metric, and the pair $(\mathbb{X},\delta)$ is called a metric space [6]. If the distance does not satisfy the strict positiveness property $(p4)$ then the space is called a pseudo-metric space, as well as if it does not hold the symmetry property $(p2)$, we may have a quasi-metric space. Nevertheless, there are some primitive spaces where these metric constraints are not necessarily satisfied. We also use the word ’dissimilarity’ when referring to this difference or distance between elements [18].
Normed and inner product spaces are special cases of metric spaces, where a metric is defined either by a norm or an inner product. An Euclidean space is a specific case of a Hilbert space. In turn, a Hilbert space is an inner product space and a particular case of a Banach space, an example of normed spaces in which a norm metric $\ell norm$ has been defined [5].
3.3 Minkowski distance
Minkowski distances or $\ell_{p}norms$ are distance measures commonly used to compute the dissimilarity of objects described by numerical attributes. However, dissimilarity can also be quantified using non-numerical attributes through other alternatives, such as the edit distance widely used for text attributes.
Let $\ell_{\infty}^{p},p\geq 1$ be a vector space of real sequences such that $\sum_{i=1}^{\infty}|x_{i}|^{p}$ with the norm given by $||x||_{p}(\sum_{i=1}^{\infty}|x_{i}|^{p})^{\frac{1}{p}}$. This norm induces the $\ell norm$ or Minkowski metric $d_{p}$ and consequently $\ell_{p}^{m}$ and $\ell_{p}^{m}=(\mathbb{R}^{m},d_{p})$ are Banach spaces.
Thus, assuming $x,y\in\mathbb{R}^{m}$, the $\ell norm$ or Minkowski distance is defined as follows:
$d_{p}(x,y)=(\sum_{i=1}^{m}|x_{i}-y_{i}|^{p})^{\frac{1}{p}},p\geq 1$ (2)
where depending on the $p$ value we can induce:
  • $d_{1}(x,y)=\sum_{i=1}^{m}|x_{i}-y_{i}|$, the Manhattan or city-block distance.
  • $d_{2}(x,y)=d_{E}(x,y)=(\sum_{i=1}^{m}|x_{i}-y_{i}|^{2})^{\frac{1}{2}}$, the Euclidean distance.
[…]
Minkowski distances with $p<1$ cannot be considered as proper metrics since they do not hold the triangle inequality. Nonetheless, several authors have shown that fractional distances can significantly improve the effectiveness of standard clustering algorithms such as $k$-means [22].
3.4 Limitations of k-means and the Euclidean distance
In unsupervised machine learning with $k$-means, the user must determine the appropriate number of clusters for a given problem, and a centroid represents each cluster. Carefully choosing the number of centroids is crucial because too many centroids would not effectively summarise the underlying data, and too few centroids would result in grouping unrelated data points. However, the locations provided to these initial centroids can greatly impact the final outcome, making it a sensitive aspect.
To address this limitation, algorithms such as $k$-means++ [27] propose an approach where the initial random centroids are spread more evenly, leading to improved results. Nevertheless, although several more powerful $k$-means implementations have tried to reduce the performance impact resulting from sub-optimal initial centroid placement, only a few have explored the potential for achieving better solutions by choosing a more appropriate distance.
Due to the nature of the $k$-means algorithm, the choice of distance function can also exert a significant impact on $k$-means outcomes, as it determines how similarity between data points is calculated and produces clusters with different shapes and separations between each other. As it has been raised in Section “Background and related work”, despite the widespread use of Euclidean distance in indexing structures and algorithms, relying on it as the default option when using the $k$-means algorithm can lead to poor results.
Previous research [22] suggests that inherent characteristics of high-dimensional spaces, such as high dispersion, outliers or concentration of points around specific values, hinder the discriminatory power of the $\ell_{p}norm$ for higher values of $p$. This situation would provide undue importance to the distant (sparse/noisy) dimensions and poorly discriminate between further and nearer neighbours. Therefore, for a given high-dimensional problem, it may be preferable to use Minkowski distances with lower values of $p$, so the $\ell_{1}$ distance metric (Manhattan distance) would be the best choice for high-dimensional applications, followed by the Euclidean distance ($\ell_{2}$), then the $\ell_{3}$ metric, up to the Chebyshev distance ($\ell_{\infty}$).
Previous algorithms for approximate similarity search, such as MASK [2], have relied on the Euclidean distance to calculate relative dissimilarities between points. This strategy follows the conventional inclination to associate distance between points with the Euclidean metric, which aligns with the distance concept commonly applied in geospatial problems. However, the potential for achieving improved results, coupled with the lack of prior empirical studies that validate this proposal and the need for learning paradigms applicable to a broader range of dissimilarity measures [5], justify our efforts to explore alternative Minkowski distances for data indexing. This endeavour aims to foster a deeper understanding of the impact of using alternative distance metrics with high-dimensional data.
Nevertheless, while it may seem theoretically plausible to employ $k$-means or its enhanced version, $k$-means++, with several distance metrics, the practical implementation reveals that the algorithm’s internal design is intertwined with the Euclidean distance. This alignment relies on the algorithm’s iterative optimisation process when it searches for centroids that minimise the total squared distances of points within their respective clusters[3]. The sum of the Squared Error (SSE), calculated as the sum of squared Euclidean distances between each data point and the centroid of its assigned cluster, underscores the algorithm’s inherent reliance on Euclidean geometry for clustering optimisation [28].
As we have already demonstrated, assuming that widely adopted mechanisms are the most suitable or the only viable option can lead to self-imposed limitations that generate sub-optimal outcomes. Instead, our objective is to steer clear of the misconception that certain implementations are the best or the only conceivable choice simply because they are the most common choices.
At this juncture, GMASK can integrate an alternative clustering algorithm that leverages the opportunities provided by diverse distance metrics for data indexing. The $k$-means algorithm defines clusters based on centroids minimising the sum of squared errors in distances from all other points in the same cluster. In contrast, alternative methods like $k$-medoids calculate clusters around the most representative data element, which is determined by minimising the sum of pairwise dissimilarities among data points the same cluster.
3.5 k-medoids as a generalisation of k-means
Several generalisations of the $k$-means algorithm have been proposed to broaden the scope of applicable data types [29], reflecting the diversity found in nature prototype points, such as subsets of objects, hyperplanes, etc.
The first paper proposes a general criterion and its generalised $k$-means method is [30]. This approach starts from a $n$ x $n$ dissimilarity matrix $(d_{\ell}t)$ for $n$ factories $O=1,...,n$ in an industrial network where $d_{\ell}t$ are the minimum road transportation costs between $\ell$ and $t$. The goal is to partition the set of factories into $k$ classes $(C_{1},...,C_{k})$ and to find a selection $Z=(z_{1},...,z_{k})$ of $k$ factories as ’supply points’ such that, when supplying all factories of the class $C_{i}$ from the supply point $z_{i}\in O$, the overall transport costs are minimised, where $d(\ell,z_{i})=d_{\ell},z_{i}$ means the dissimilarity between the factory (object) $\ell$ and the factory (supply point) $z_{i}\in O$. So the family $P$ of admissible ’prototypes’ consists of all singletons from $O$ and the class prototypes, meaning ’most cheapest supply point’ in $C_{i}$, are determined by:
$Q(C_{i},z):=\sum_{\ell\in C_{i}}d(\ell,z)\to\min_{z\in C_{i}}$ (3)
Many authors, including, among others, Diday [31, 32, 33], Bock [34] or Govaert [35], have followed the generalised clustering approach in various settings and numerous variations, obtaining a plethora of generalised $k$-means algorithms.
Kaufmann and Rousseeuw [36] first propose the popular Partitioning Around Medoids (PAM) algorithm, also referred to as $k$-medoids. In $k$-medoids, data is modelled by using $k$ representative objects $m_{i}$ called medoids, that serve as prototypes for the clusters (formerly called classes) which are obtained by using the absolute error criterion (’total deviation’, TD). Since the TD is defined as the sum of dissimilarities of each point $x_{j}\in C_{i}$ to the medoid $m_{i}$ of its cluster, the medoid of a cluster $C_{i}$ is in fact the object with the smallest sum of dissimilarities (or equivalently, the smallest average) to all other objects in the cluster.
$medoid(C_{i}):=argmin_{x_{i}\in C}\sum_{{x_{j}\in C}}d(x_{i},x_{j})$ (4)
This method is suitable for non-Euclidean distances and a wide range of data since it does not force the dissimilarity measure to adhere strictly to metric properties, and only requires a proximity measure between any pair of objects. In addition, while $k$-means defines a prototype in terms of a centroid, which is usually the mean of a group of points that may not correspond to an element in the dataset, $k$-medoids defines a prototype in terms of a medoid, an actual data point identified as the most representative one for a group of points.
Regarding $k$-medoid implementations, Program PAM [36] consists of two algorithms, BUILD to choose an initial clustering, and SWAP to improve the clustering towards a local optimum, since finding the global optimum of the $k$-medoids problem is, unfortunately, NP-hard. The algorithms require a dissimilarity matrix, which requires $O(n^{2})$ memory and at least $O(n^{2}d)$ time to be computed. On top of this, the process of updating medoids involves evaluating the cost of swapping non-medoid points to ensure that the sum of pairwise dissimilarities within clusters (rather than the sum of squared Euclidean distances) is minimised, making this method more robust to noise and outliers compared to the alternatives based on $k$-means. Recently, FasterPAM [37] introduced a speedup for larger $k$, achieving an $O(k)$-fold acceleration in the second SWAP phase of the algorithm, but still providing the same results as the original PAM algorithm.
Originally FasterPAM was implemented in Java and published within the open-source library ELKI [38]. However, to integrate it with GMASK implementation, developed entirely in Python, we will employ the fast $k$-medoids Python implementation [39] version provided by the $kmedoids$ package  [^1] https://pypi.org/project/kmedoids. FasterPAM let us upgrade our method by making it compatible with any distance function and a wider range of problems.
Experimental results verify the applicability of GMASK with real datasets, improving the performance of alternative algorithms and providing better results for approximate similarity search queries.
3.6 Performance evaluation
The main advantage of GMASK is providing a fast response to queries at the expense of returning approximate results. Thus, a straightforward strategy to assess the performance of this algorithm is to undertake an exhaustive search of all elements in a given dataset $X$. Then, the proportion of data points that were correctly retrieved can be determined. This is the common approach to all evaluation experiments developed in this study.
A comprehensive series of experiments are conducted across datasets of diverse types, sizes, dimensionality, and sparsity to evaluate the performance of GMASK, which introduces $k$-medoids for the bottom-top construction of the index tree. This novelty enables its parametrisation with several $\ell_{p}$ norms as outlined in “k-medoids as a generalisation of k-means”. These results will be compared with those obtained using other state-of-the-art algorithms. As mentioned above, our objective is to understand better the impact of utilising distances other than the Euclidean metric. Through these experiments, we aim to demonstrate that data indexing using non-Euclidean distances is not only feasible but also has the potential to yield improved outcomes.
3.7 Benchmark algorithms for similarity search
To evaluate our thesis, we compare GMASK performance in approximate nearest neighbour search with respect to alternative MAM.
3.7.1 KD-Tree
The KD-Tree (K-Dimensional Tree) [40] is an exact similarity search method used to find the nearest neighbours to a given query point. It constructs a binary tree structure by recursively dividing the data based on the maximum spread along different dimensions and the search algorithm efficiently narrows down the search space by traversing the tree. In low-dimensional spaces, KD-trees can achieve sublinear search times. However, as the dimensionality increases, the performance of the KD-tree can degrade due to the Curse of Dimensionality phenomenon.
As it performs an exact search, the solutions provided by this method are used as a baseline to evaluate the performance of other approximate algorithms. We shall compare the k-nearest neighbours obtained by each method against those obtained using the KD-Tree to assess their accuracy.
3.7.2 FLANN
FLANN (Fast Library for Approximate Nearest Neighbours) [11] is an efficient method widely employed for approximate similarity search in large-scale and high-dimensional datasets, extensively used in machine learning, computer vision, and data mining. FLANN employs techniques such as random projection and hierarchical subdivision to construct an index structure that enables faster search operations. It offers several index types, including KD-tree, $k$-means tree, and hierarchical clustering tree, which are suitable for diverse data types and query requirements. While FLANN has implementations in several programming languages, we use the Python bindings provided by the pyflann  [^2] https://github.com/primetang/pyflann library.
3.7.3 PyNNDescent
The PyNNDescent algorithm is an approximate similarity search method [41] implemented in Python  [^3] https://pynndescent.readthedocs.io/en/latest/. It is based on the NNDescent algorithm and used for efficient nearest neighbour search in large-scale datasets. PyNNDescent constructs a graph-based index by connecting each data point to its approximate nearest neighbours. It iteratively improves the graph structure through neighbour descent steps, employing techniques such as random initialisation, multi-tree searching, and pruning. The algorithm also provides parameters to control the trade-off between search accuracy and computational cost. PyNNDescent is particularly useful for high-dimensional datasets and is widely used in machine learning, data mining, and information retrieval tasks.
3.8 Datasets
Several experiments have been conducted to evaluate the performance of GMASK, using four datasets diverse in nature and public access. Table 1 briefly summarises their main features.
3.8.1 Municipalities
The Municipalities dataset is a two-dimensional geospatial dataset that contains the coordinates (latitude and longitude) of each of the 8,130 municipalities that make up the Spanish territory, as depicted in Figure 3. It has been extracted from one of the tables provided in the ’2020 Resumen Descriptivo del Nomenclátor de Municipios y Entidades de Población’ (Descriptive Summary of the Nomenclature of Municipalities and Population Entities), called MUNICIPIOS [^4] https://ign.es/web/rcc-nomenclator-nacional. This report is compiled annually by the Instituto Geográfico Nacional (National Geographic Institute), a public institution under the Spanish Ministry of Transport and Sustainable Mobility. It includes updated geographic and administrative data on all Spanish municipalities, such as their name, province, geographic location coordinates, perimeter or population, among others.
Figure 3: Visualisation for dataset Municipalities
Although this is a two-dimensional dataset with a low dispersion, it is an exceptionally intriguing case due to its intuitive representation and the presence of outliers, such as the Canary Islands and the Balearic Islands. The location of these islands deviates from the rest of the points significantly and poses a challenge to any similarity search method applied to this dataset.
3.8.2 MNIST
The MNIST_784 dataset  [^5] http://yann.lecun.com/exdb/mnist/ is a widely used machine learning benchmark dataset whose simplicity makes it an ideal starting point for algorithm development and evaluation. Firstly, it consists of 60,000 training images and 10,000 test images, a relatively small size that makes possible to carry out computationally efficient experimentation and prototyping. Secondly, it comprises grayscale images of handwritten digits ranging from 0 to 9, making it visually straightforward and easy to work with. Moreover, each image is 28x28 pixels in size, resulting in a total of 784 dimensions when flattened, allowing researchers to explore algorithms in the context of high-dimensional data. Many well-established techniques and algorithms have been evaluated on MNIST [42, 43], making it a reference point for evaluating the performance of approximate similarity search methods [16, 44].
3.8.3 GLOVE
The third experiment involves the GLOVE_100 dataset [^6] https://nlp.stanford.edu/data/glove.6B.zip, which contains one million pre-trained Global Vectors for Word Representation (GloVe) embeddings that capture semantic relationships between words. These 100-dimensional embedding vectors were pre-trained on the combined Wikipedia 2014 + Gigaword 5th Edition corpora (6B tokens, 400K vocabulary) and cover a wide range of English words, enabling comprehensive assessments of approximate similarity search methods across diverse vocabulary and word relationships. Given their widespread adoption as a benchmark in the natural language processing community, the use of GloVe embeddings as a standardised dataset provides valuable insights into the method effectiveness in retrieving similar items based on their semantic similarity [16].
3.8.4 NYtimes
The NYtimes dataset, part of the Bags of Words dataset from the UCI repository [^7] https://archive.ics.uci.edu/dataset/164/bag+of+words, comprises a collection of New York Times news articles represented as a bag of words. Each document in the dataset is associated with a set of word occurrences, where the dimensions represent unique words extracted from the articles. The dataset is organised as a document–word matrix, where each row corresponds to a document and each column corresponds to a word. The values in the matrix indicate the frequency of each word occurring in the respective document. Preprocessing steps include tokenization, removal of stopwords, and vocabulary truncation, with only words occurring more than ten times retained.
The reduced version of the NYtimes employed on this work was generated by building tf-idf descriptors from the bag-of-words version, and embedding them into a lower dimensional space using the Johnson-Lindenstrauss Transform [45].
Dataset Label n Dimensionality High Sparsity Data Type
Municipalities $\mathcal{X}_{muni}$ 8,130 2 No Geospatial
MNIST $\mathcal{X}_{MNIST}$ 69,000 784 Yes Image
GLOVE $\mathcal{X}_{GLOVE}$ 1,000,000 100 No Text
NYtimes $\mathcal{X}_{NYtimes}$ 290,000 256 No Text
Table 1: Descriptive traits characterising the datasets included in experiments.
4 Experimental results
In order to show the versatility of GMASK, we present the analysis of four different datasets: municipalities, MNIST, GLOVE and NYtimes. This study consists of a neighbours search applying Euclidean distance, Manhattan distance and Chebyshev distance. The main objective is to measure the accuracy of the GMASK algorithm with different distances and to observe if any distance is more adequate for some dataset, depending on its characteristics. For each dataset, we compare the results obtained by GMASK with two other algorithms: PYNN and FLANN. Both PYNN and GMASK are able to work with these three distances, however, FLANN does not work with Chebyshev distance. Since MASK is a specific case of GMASK only able to work with Euclidean distance, its inclusion in these experiments has not been considered.
Our approach to evaluating GMASK performance is conceived to closely emulate real-world conditions. The datasets used in this work are not artificially generated, and no preprocessing has been applied. The transformation and cleaning of the collected data is limited to dividing the original dataset into two distinct subgroups, the training and testing datasets. We also check that identical elements do not appear in both subgroups, thereby intensifying the complexity in the search for $k$ approximate neighbours.
All experiments have been performed on a server equipped with 2 AMD EPYC 7451 microprocessors (24 cores/48 threads, 2.30 GHz, 64 MB L3 cache), 128 GB of DDR4 RAM and an Intel D3-S4510 SSD (capacity 480 GB) for secondary storage. The experimental results are presented in the following sections.
4.1 Municipalities
The municipalities dataset is an interesting case due to the geographical distribution of points. In Figure 3, we can identify three main groups of points: the largest region corresponds to the peninsular territory of Spain, whereas the group of points located at the right of the map corresponds to the Balearic Islands and the group of points located at the bottom left corner of the map corresponds to the Canary Islands. The last two groups of points increase the difficulty of the analysis because they are far from the rest of the points, so we can consider them as outliers. Despite this dataset containing additional information, we only consider two variables: latitude and longitude.
Figure 4 shows the recall comparison between FLANN, PYNN and GMASK. We use a heatmap to represent the recall value obtained for each algorithm with each distance in the search of 5, 10 and 15 neighbours. The recall obtained in this case for each method is very similar, and all of them depict good recall values, over 84%. It is noticeable that FLANN recall decreases by about 10% with the number of neighbours. Regardless of the distance used and the number of neighbours intended to search, GMASK is the only method that provides a 100% recall in all tests.
Figure 4: Heatmap comparing the Recall (percentage of true nearest neighbours found) obtained by each algorithm and distance in approximate nearest neighbour search, for different number of neighbours, with dataset $\mathcal{X}_{muni}$.
4.2 MNIST
Given its high dimensionality and sparsity, MNIST has become a widespread benchmark in similarity search research. The following experiment evaluates the recall obtained in searching neighbours in this dataset, applying Manhattan, Euclidean and Chebyshev distance with FLANN, PYNN and GMASK methods. Figure 5 presents the results of this analysis.
From this comparison, we can deduce that the results are highly comparable in the recall when applying PYNN and GMASK methods with Euclidean and Manhattan distances, respectively. Therefore, the combination of both distances and these methods suits this case study. However, this is not the case for Chebyshev distance since the recall is lower than 70%. There is also a meaningful reduction in FLANN recall when the number of neighbours increases. Regardless of the distance, GMASK yields a higher average recall than other methods.
Figure 5: Heatmap comparing the Recall (percentage of true nearest neighbours found) obtained by each algorithm and distance in approximate nearest neighbour search, for different number of neighbours, with dataset $\mathcal{X}_{MNIST}$.
4.3 GLOVE
Hereafter, the third experiment involves the analysis of the GLOVE dataset, a high dimensional datasets that contains the largest number of elements among the datasets included in these experiments.
Figure 6 depicts the recall results obtained for each method and distance. Once again, the luck of efficiency is evident in FLANN´s results. The best recall obtained with FLANN is 12% when searching for 5 neighbours applying Euclidean distance. PYNN method provides good accuracy when Manhattan distance is used. If we focus on PYNN method and compare the recall for the three distances, we can see that Manhattan distance provides the best results and Chebyshev distance provides the worst recalls. However, in the case of GMASK, the recall is 100% per each distance and with independence of the number of neighbours searched.
Figure 6: Heatmap comparing the Recall (percentage of true nearest neighbours found) obtained by each algorithm and distance in approximate nearest neighbour search, for different number of neighbours, with dataset $\mathcal{X}_{GLOVE}$.
4.4 NYtimes
Finally, the last experiment replicates the same analysis using a dataset characterised by high dimensionality, featuring a data-to-dimensionality ratio lower than that of GLOVE but higher than MNIST: the NYtimes dataset.
Figure 7 illustrates the recall accomplished by FLANN, PYNN and GMASK. There is a noticeable difference between GMASK and the other methods, since the recall values for FLANN and PYNN are below 50%. However, GMASK consistently outperforms the other methods regardless of the distance chosen. Once again, FLANN exhibits the poorest performance.
Therefore, while the most remarkable insight obtained from this experiment is the superior performance of the Chebyshev distance metric compared to others, it is crucial to highlight that, in this case, GMASK nearly doubles the recall provided by PYNN, achieving an 82% recall when searching for the 5 nearest neighbours.
Through this experiment, we empirically demonstrate what was previously stated theoretically in earlier sections of this manuscript: adopting the Euclidean distance as the default option in indexing structures and algorithms is a self-imposed restriction that can lead to sub-optimal results.
Figure 7: Heatmap comparing the Recall (percentage of true nearest neighbours found) obtained by each algorithm and distance in approximate nearest neighbour search, for different number of neighbours, with dataset $\mathcal{X}_{NYtimes}$.
5 Discussion
Given the notable prevalence of Euclidean distance in similarity searches in the previous literature, as well as the limitations detected in the widely adopted $k$-means algorithm, which is strongly linked to this metric, there is a growing need to explore the use of alternative metrics when addressing these types of problems. Stemming from this need, GMASK makes the accommodation of arbitrary distances feasible by integrating the $k$-medoids algorithm as a clustering algorithm, allowing us to leverage the advantages derived from using several distance metrics on high-dimensional datasets.
It is crucial to emphasise that the algorithm’s generalisation, achieved through the integration of $k$-medoids, significantly broadens the range of applicable distances while preserving the structural integrity and properties of the original algorithm. Thus, GMASK maintains crucial characteristics, including the desirable scalability property facilitated by the bottom-top construction of the search tree, retaining a competitive edge over other state-of-the-art algorithms that require centralising all data on a single server.
As elaborated in Section “Experimental results”, our findings demonstrate the efficiency of this new approach, thereby supporting the hypothesis that alternative distances beyond the Euclidean metric may yield superior results in solving nearest-neighbour queries, contingent upon the specific characteristics of the dataset. Therefore, it becomes evident that the outcomes achieved using the proposed method surpass those obtained through two other widely adopted state-of-the-art algorithms, which are distinct from each other and significantly differ from GMASK.
These findings would also be in line with previous publications [22] suggesting that, for a given high-dimensional problem, using an $\ell_{p}$ norm with lower values of $p$ may be advantageous in assessing dissimilarity between query points. Thus, while Aggarwal carries out a theoretical approach using synthetic data, our research delves into this hypothesis by generalising an algorithm that has been tested over a range of real-world datasets of diverse nature. However, further empirical research is required to validate this observed behaviour more precisely.
GMASK provides an avenue for future research to integrate distance functions beyond Minkowski, thus providing many opportunities to extend this research line. For instance, it would be interesting to explore the suitability of distance functions for specific purposes, such as cosine similarity for textual data or the Jaccard distance for categorical data, to address particular problems.
Furthermore, although GMASK has yielded notably satisfactory experimental results, there still exists potential for enhancing this implementation to achieve further optimisation in computational methods, thereby increasing overall performance. To this end, the proposal suggests exploring the incorporation of a just-in-time (JIT) compiler for Python, such as Numba  [^8] https://numba.pydata.org—employed by other algorithms such as PYNNDescent—, or considering the adaptation and port of GMASK to a language renowned for its strong performance in building scalable and concurrent systems like Go  [^9] https://go.dev.
6 Conclusions and future work
Approximate similarity search is essential for many application areas where large and heterogeneous datasets arise, often exhibiting high dimensionality and sparsity. In many cases, these datasets must be stored in distributed systems to become manageable, imposing additional restrictions that rule out algorithms conceived for cases where data must reside on a single machine.
This paper presents the GMASK algorithm for approximate similarity search in distributed systems that can accept arbitrary distance functions. One of the main novelties is adopting the $k$-medoids (PAM) algorithm, which is not bounded to the Euclidean distance, to achieve efficient solutions. Furthermore, using FastPAM, a highly efficient implementation of the PAM algorithm to solve the $k$-medoids problem, GMASK can improve its efficiency with respect to other previous algorithms.
Experiments conducted on real-world datasets with different properties confirm that GMASK, based on $k$-medoids, performs better in solving nearest-neighbour queries than other frequently used algorithms, using other Minkowski dissimilarity functions different from the Euclidean distance. These results illustrate that indexing data with non-Euclidean distances is not only viable but also holds the potential to enhance search outcomes substantially. Overlooking alternative distances would represent a voluntary limitation that may hinder us from achieving the optimal results.
GMASK lays out new interesting routes for further research. Once the requirement of using the Euclidean distance has been lifted, additional studies should be conducted to evaluate GMASK performance in other settings where alternative distances can be applied, such as pseudo-metrics (Minkowski distances with $p<1$) or non-metric dissimilarities. Besides, using arbitrary distances will enable us to consider datasets with qualitative features exclusively or a combination of quantitative and qualitative features. Many state-of-the-art algorithms and standard dissimilarity measures cannot be directly applied in these cases since they are incompatible with categorical features.

### v2

@@
-Using $k$-medoids for distributed approximate similarity search with arbitrary distances
+A general framework for distributed approximate similarity search with arbitrary distances
 Abstract
-This paper presents GMASK, a general algorithm for distributed approximate similarity search that accepts any arbitrary distance function.
-GMASK requires a clustering algorithm that induces Voronoi regions in a dataset and returns a representative element for each region.
-Then, it creates a multilevel indexing structure suitable for large datasets with high dimensionality and sparsity, usually stored in distributed systems.
-Many similarity search algorithms rely on $k$-means, typically associated with the Euclidean distance, which is inappropriate for specific problems.
-Instead, in this work we implement GMASK using $k$-medoids to make it compatible with any distance and a wider range of problems.
-Experimental results verify the applicability of this method with real datasets, improving the performance of alternative algorithms for approximate similarity search.
-In addition, results confirm existing intuitions regarding the advantages of using certain instances of the Minkowski distance in high-dimensional datasets.
-Keywords Approximate similarity search  $\cdot$ k-medoids  $\cdot$ arbitrary distance function
+Similarity search is a central problem in domains such as information management and retrieval or data analysis.
+Many similarity search algorithms are designed or specifically adapted to metric distances.
+Thus, they are unsuitable for alternatives like the cosine distance, which has become quite common, for example, with embeddings and in text mining.
+This paper presents GDASC (General Distributed Approximate Similarity search with Clustering), a general framework for distributed approximate similarity search that accepts arbitrary distances.
+This framework can build a multilevel index structure, by selecting a clustering algorithm, the number of prototypes in each cluster and any arbitrary distance function.
+As a result, this framework effectively overcomes the limitation of using metric distances and can address situations involving cosine similarity or other non-standard similarity measures.
+Experimental results using $k$-medoids clustering in GDASC with real datasets confirm the applicability of this approach for approximate similarity search, improving the performance of extant algorithms for this purpose.
+Keywords Data indexing  $\cdot$ Approximate similarity search  $\cdot$ Arbitrary distance  $\cdot$ Clustering  $\cdot$ Unsupervised learning
 1 Introduction
-The goal of similarity search is to find elements from a dataset, represented by feature vectors in a multidimensional space, similar to a given query object according to a particular similarity or dissimilarity measure.
+The goal of similarity search is to find elements from a dataset, represented by feature vectors in a multidimensional space, similar to a given query object according to a similarity or dissimilarity measure.
 Numerous data indexing methods, collectively known as access methods, have been proposed to speed up this task.
-Nevertheless, the results remain strongly conditioned by the problems inherent to the representation of datasets in high-dimensionality and high-sparsity spaces.
-These algorithms are frequently conceived to work with data having spatial components or represented in a vector space with few dimensions.
-However, they become unusable in high-dimensional spaces, where many features characterise each element due to the Curse of Dimensionality [1].
-As a result, the distance between elements tends to become increasingly uniform, making it challenging to differentiate between them based solely on their distances or when the number of dimensions significantly exceeds the quantity of data points.
-Moreover, several studies indicate that in datasets with over 20 dimensions, exact similarity search methods rarely outperform exhaustive scanning.
-Owing to this constraint, novel approaches have emerged in recent years to swiftly identify enough related items within a reasonable time frame compared to linear scanning, often referred to as approximate similarity search.
-One of the newest approaches is MASK (Multilevel Approximate Similarity search with k-means) [2], a novel indexing method that involves an unconventional application of the $k$-means partitioning algorithm [3, 4] to create a multilevel index structure for approximate similarity search.
-Traditionally, the input to $k$-means is always a fixed number of $k$ prototype points or centroids.
-Then, following an iterative process, the dataset is divided into $k$ groups (Voronoi cells), assigning each point to its closest centroid according to a given distance function, so that each centroid represents all of the elements of that cluster.
-In contrast, leveraging current powerful computing capabilities, $k$-means can be forced to generate a high number of centroids that inundate the dataset.
-Then, these centroids become a new dataset to be grouped at the upper level, repeating this process until the complete construction of the search tree is achieved.
-This bottom-up creation of the multilevel index enables the distribution of this task among different computing nodes, which also store distinct data partitions.
-Although it employs an innovative approach that capitalises on the benefits offered by contemporary computing systems and distributed platforms, the performance of this indexing method remains significantly impacted by the inherent constraints of the $k$-means algorithm.
-The choice of the distance function significantly impacts the outcome of $k$-means clustering, as it determines how the similarity between data points is calculated and can lead to different outcomes in terms of the shape and separation of the clusters.
-Until now, experiments conducted with $k$-means have relied on the Euclidean distance to calculate relative distances between points, following the conventional inclination to associate the concept of distance with that used in geospatial contexts.
-Nonetheless, relying on the Euclidean distance as the default option in indexing structures and algorithms when working with high-dimensional data, despite its widespread use, can lead to sub-optimal results.
-This is because, in high-dimensional spaces, the discriminating power of the $L_{p}$ norm decreases for higher values of $p$ due to their specific features (such as dispersion, outliers, and concentration of points around certain values), leading to poor differentiation between proximate and distant elements.
-This paper introduces GMASK (General Multilevel Approximate Similarity Search with $k$-medoids), a novel algorithm that facilitates the incorporation of arbitrary distances by employing data partitioning algorithms that induce Voronoi regions in a dataset and yield a representative element, exemplified by $k$-medoids.
-The simple introduction of $k$-medoids turns GMASK into a versatile algorithm capable of adapting to the characteristics of any dataset (high dispersion, high dimensionality, outliers, etc.) and seek for the optimal results.
-This approach allows using multiple $\ell_{p}$ norms or even other distance functions, making it compatible with a broader spectrum of problems.
+Nevertheless, their performance remains strongly conditioned by complications inherent to dataset representation in high-dimensional and sparse spaces.
+These algorithms are often designed to handle data with spatial components or to operate within a low-dimensional vector space.
+However, they become unusable in high-dimensional spaces, where each element is characterised by numerous features, due to the Curse of Dimensionality Böhm et al.
+[1].
+In such spaces, the distance between elements becomes increasingly uniform, making it challenging to differentiate between them.
+Analogous issues arise when the number of dimensions significantly exceeds the number of data elements.
+Moreover, several studies [2] indicate that in spaces exceeding 10 dimensions exact similarity search methods rarely outperform exhaustive scanning.
+To address these challenges, researchers have explored various techniques to reduce the dimensionality or transform the feature space to make high-dimensional data easier to manage.
+A common approach is using dimensionality reduction techniques such as Principal Component Analysis (PCA) and t-Distributed Stochastic Neighbour Embedding (t-SNE) [3].
+These techniques project data into lower-dimensional spaces while preserving as much variance or similarity structure as possible.
+However, these methods are not sufficient, especially if the intrinsic dimensionality remain high after the conversion.
+At the same time, approximate similarity search methods have become increasingly important since they attain a practical balance between accuracy and computational efficiency.
+Methods such as Locality-Sensitive Hashing (LSH) [4] and Random Projection Trees [5] aim to quickly identify nearest neighbours by mapping high-dimensional data into a space where proximity is preserved with high probability.
+These methods have proven effective in reducing the computational burden associated with high-dimensional similarity searches, although they introduce some trade-offs in terms of accuracy.
+Furthermore, hierarchical and partitioning-based methods have also been proposed for multidimensional data search.
+Methods such as the Ball-Tree [6] or the KD-Tree [7] divide the data space into hierarchical structures, accelerating queries by reducing the number of required distance calculations.
+However, the performance of these structures may degrade as dimensionality increases and, hence, more robust indexing structures must be used for high-dimensional spaces.
+Recent advances have focused on using machine learning models to improve similarity search in high-dimensional spaces.
+One such approach is deep learning embeddings, where neural networks are trained to map high-dimensional data into lower-dimensional vectors that preserve similarity relationships, which have offered significant improvements in different applications [8].
+Although effective, these methods also require significant computing resources and extensive training data to achieve optimal performance.
+Finally, the development of similarity search methods highlights a permanent trade-off between computational efficiency and accuracy in addressing high-dimensional data.
+Future research is likely to continue focusing on hybrid approaches that combine traditional indexing methods with advanced machine learning techniques to push the boundaries of what can be achieved in high-dimensional similarity search.
+In this regard, it is also worth noticing the emergence of new datasets demanding non-metric distances.
+In practice, this additional prerequisite discards many existing methods for exact or approximate similarity search that has been conceived specifically for metric spaces.
+To overcome these limitations, this paper introduces GDASC, a novel framework that allows arbitrary distances by employing different clustering algorithms that can assign a representative element to each cluster.
+These properties turn GDASC into a versatile tool that can be adapted to the specific traits of any dataset (high dispersion, high dimensionality, outliers, etc.), seeking for optimal results.
+This approach can utilise multiple $\ell_{p}$ norms besides the Euclidean distance or even other distance functions, such as cosine distance, making it suitable for a broader spectrum of problems.
 To comprehensively evaluate its performance, an extensive series of experiments will be conducted across datasets varying in type, size, dimensionality, and sparsity.
-This paper adheres to the following structure.
-Section “Background and related work” reviews previous research related to similarity search.
-Section “Methodology” describes GMASK, emphasising the potential for enhancing the algorithm’s performance by exploring the application of other distance functions beyond the Euclidean one, like other Minkowski distances.
-The findings of four empirical experiments carried out over real-world datasets with quite diverse attributes are presented in Section “Experimental results”.
-These tests compare GMASK performance with other well-known algorithms, such as FLANN and PyNNdescent.
-Finally, Section “Conclusions and future work” recaps the main conclusions of this work and suggests promising lines for future work.
+The paper adheres to the following structure.
+Section 2 reviews previous research related to similarity search.
+Section 3 describes GDASC, emphasising the potential for enhancing the algorithm’s performance by exploring the application of other distance functions beyond the Euclidean one.
+The findings of four empirical experiments carried out over real-world datasets with quite diverse attributes are presented in Section 4.
+These tests compare GDASC performance with other well-known algorithms, such as FLANN and PyNN.
+The implications of using arbitrary distances and different clustering algorithms in similarity searches are discussed in Section 5.
+Finally, Section 6 recaps the main conclusions of this work and suggests promising lines for future work.
 1.1 Contributions
-The potential for achieving improved results, coupled with the lack of prior empirical studies that validate this proposal and the need for learning paradigms applicable to a wider range of dissimilarity measures [5], justifies our efforts to explore alternative Minkowski distances for data indexing.
-This endeavour is aimed at challenging the misconception we have identified and fostering a deeper understanding of the significant impact of using alternative distance metrics when working with high-dimensional data.
-  • GMASK allows the use of different clustering algorithms for the bottom-top construction of the index tree, such as $k$-medoids, a broader form of the $k$-means algorithm that offers parameterisation with several distance metrics, in particular Minkowski distances.
-  • Evaluating the performance of different $L_{p}$ norms or Minkowski distances for approximate $k$-nearest neighbour search.
-This assessment involves testing GMASK against other state-of-the-art methods to illustrate that the selection of a suitable distance function, while considering the features of the target dataset, has a significant impact on performance when searching for reasonably similar regions.
+The potential for achieving improved results, coupled with the lack of prior empirical studies that validate this proposal and the need for learning paradigms applicable to a wider range of dissimilarity measures [9], justifies our efforts to explore alternative distance metrics for data indexing.
+This endeavour is aimed at challenging the misconception we have identified and fostering a deeper understanding of the significant impact of using alternative distance metrics when working with high-dimensional data [10].
+  • Proposing GDASC, a versatile framework for data indexing that supports the use of arbitrary distances by employing different clustering algorithms for the bottom-up construction of the index tree, such as $k$-medoids.
+  • Evaluating the performance of multiple $\ell_{p}$ norms or Minkowski distances as well as cosine distance for approximate $k$-nearest neighbour search.
+This assessment involves testing GDASC against other state-of-the-art methods to illustrate that the selection of a suitable distance function, while considering the features of the target dataset, significantly impacts performance when searching for reasonably similar regions.
 2 Background and related work
@@
 2.1 Similarity search
-Similarity search, often referred to as proximity search [6, 7, 8], is a cornerstone for applications in many different fields, such as databases [9], information retrieval [10], computer vision [11] and bioinformatics [12, 13], among others.
-It aims to find elements that are similar (close) to a given query object, according to some measure of similarity or, conversely, dissimilarity.
-Thus, given a query object $q\in\mathbb{X}$, the similarity search tackles some of the following scenarios, represented in Figure 1 [7]:
-  • Point query: Finding elements within $\mathbb{X}$ that have exactly the same feature values as $q$.
-  • Range query: Identifying a subset of elements $\{o_{i}\}\in\mathbb{X}$ whose feature values do not exceed a given similarity threshold $r$ around $q$ such that $\forall\{o_{i}\},\delta(o_{i},q)\leq r$.
-  • Nearest-neighbour query: Finding elements in $\mathbb{X}$ whose features are the closest to the feature vector representing $q$.
-In this case, the goal might be to retrieve the single element (set of $k$ elements) closest to $q$, known as the nearest neighbour ($k$-nearest neighbours) and denoted by $NN_{q}$ ($k-NN_{q}$).
-Identifying the $k$-nearest neighbours of a given object is one of the most widely used classification methods in machine learning.
+Similarity search [11, 12, 13] is an essential technique underpinning applications in a wide range of fields, including but not limited to information retrieval [14], computer vision [15], machine learning [16] and bioinformatics [17, 18].
+It is an information retrieval technique involving a set of objects characterised by feature vectors and represented as points in a multidimensional space.
+The goal is to find the closest (most similar) object to a given query, typically determined by some similarity or dissimilarity measure, often expressed as a distance between two objects.
+Let $\mathcal{X}$ denote a set of objects represented by feature vectors in a multidimensional space and $\delta:\mathcal{X}\times\mathcal{X}\mapsto\mathbb{R}$ is a function that assigns a real number $\delta(a,b)$ to each pair of objects $a,b\in\mathcal{X}$, indicating the distance between them.
+Thus, given a query object $q\in\mathbb{X}$, similarity search can deal with different scenarios, as depicted in Figure 1 [12]:
+  • Range query: Finding all objects of $\mathbb{X}$ within a specified threshold around $q$. Formally, retrieving $S=\{o_{i}\mid\delta(o_{i},q)\leq r,\forall o_{i}\in\mathbb{X}\}$.
+  • Nearest-neighbour query: Finding the set of objects in $\mathbb{X}$ that are closest to $q$, that is, $S=\{o_{i}\in\mathbb{X}\mid\delta(o_{i},q)\leq\delta(o_{n},q),\forall o_{n}\in\mathbb{X}\}$.
+The $k$-nearest neighbours method is recognised as the primary approach in this type of search, involving the identification of the $k$ closest objects to $q$ within $\mathbb{X}$.
+Identifying the $k$-nearest neighbours is a fundamental method in machine learning classification.
 Figure 1: Conceptual representation of several types of query searches.
-Many techniques for data indexing, collectively known as access methods, can speed up these queries.
-These methods create an index structure that partitions the domain $\mathbb{X}$ into different regions based on a distance function.
-Then, we can leverage these regions to solve similarity queries.
-There is a wide range of access methods, including exact approaches, approximate solutions and variants tailored for distributed computing systems.
+Many approaches to accelerate similarity queries have been proposed, most of them based on access methods.
+These methods use efficient indexing structures that divide the representation space into regions defined by a distance function.
+Access methods include a wide range of techniques, ranging from exact methods to approximate solutions, and adaptations designed for distributed computing systems.
 2.2 Algorithms for similarity search
 Exact similarity search methods identify the completely accurate result set for a query by using the absolute positions of objects to group elements and search for similar instances.
-These methods are categorised as Point Access Methods (PAM) for point-based elements and Spatial Access Methods (SAM) for elements with spatial dimensions, such as lines, polygons or higher-dimensional polyhedra.
+These methods are categorised as Point Access Methods (PAM) for point-based elements, and Spatial Access Methods (SAM) for elements with spatial dimensions, such as lines, polygons or higher-dimensional polyhedra.
 The general acronym SAM is commonly used in the research literature to refer to both categories of indexing strategies in vector spaces [14].
-These approaches are most effective with elements represented in a relatively small number of dimensions, making Curse of Dimensionality one of the biggest challenges in data science today.
-Metric Access Methods (MAM) provide a general framework for similarity search problems within metric spaces [6, 8], exploiting the triangle inequality to partition the metric space into subspaces, effectively filtering out portions of the dataset that cannot contain valid results.
+These approaches are most effective with elements represented in a relatively small number of dimensions, making the Curse of Dimensionality in high-dimensional spaces one of the biggest challenges in data science today.
+Metric Access Methods (MAM) provide a general framework for similarity search problems within metric spaces [11, 13], exploiting specific properties like the triangle inequality to partition the metric space into subspaces, effectively filtering out portions of the dataset that cannot contain valid results for a query.
 On the one hand, MAM exact search can be expensive in terms of computational time and updating the index structure due to dynamic data.
@@
 There are many MAM for queries in metric spaces, some of which use clustering algorithms, such as $k$-means [19] or $k$-medoids [20], to recursively partition the entire set of elements creating a hierarchical structure (tree) of nested clusters.
-Approximate indexing methods are solutions that strike a balance between accuracy and speed, making them suitable for both vector and metric spaces [15].
-Although these methods may introduce some errors into the results, they also provide a faster response time by calculating fewer distances.
-As a result, these methods prove feasible in scenarios where exact methods cannot provide a fast response, such as in high-dimensional problems or large datasets, where exact methods default to a full scan.
-Some of the most widely used exact algorithms are compared and summarised in [16], which also introduces a comprehensive taxonomy to classify them according to relevant characteristics.
-Rather than computing the exact result set for a query, approximate similarity search algorithms adopt various strategies (e.g., modifying the representation space, reducing the number of pair distances to compute) to identify a reduced set of candidate items that are likely to be the most similar.
-FastMap, introduced in [17], is one of the first algorithms to be developed for approximate similarity search in metric spaces.
-It is based on a method to obtain a fast projection of elements in the original dataset into a user-defined $k$-dimensional space, preserving the distances between pairs of elements in order to achieve an approximate response through dimensionality reduction.
-In this category of algorithms, MASK [2] emerged as a novel indexing approach characterised by a multilevel design tailored for approximate similarity search.
-It features an innovative utilisation of the $k$-means partitioning algorithm to quickly reduce the number of checked regions when searching for results.
+Approximate indexing methods strike a balance between accuracy and speed, making them applicable to both vector and metric spaces [21].
+Despite potential inaccuracies in the results, they provide a faster response by minimising distance calculations.
+These methods are particularly useful in contexts where exact approaches are hampered by the computational burden of exhaustive scans, especially in high-dimensional datasets or scenarios involving large amounts of data
+Rather than computing the exact result set for a query, approximate similarity search algorithms use various strategies (e.g., transforming the feature space, reducing the number of pairwise distances computed) to identify a reduced set of candidate items that are likely to be highly similar.
+Some of the most commonly used approximate algorithms for speed up similarity search are tree-based methods [22], hashing-based methods [23] and graph-based methods [24].
+Tree-based methods divide the high-dimensional space hierarchically, grouping similar objects in the same partition as leaf nodes.
+In addition, objects called representatives are often used to represent each partition.
+Thus, each node stores the representatives, the objects in the covered partition, and their distances to the representatives.
+When a search query is performed, the query object is first compared to the representatives of each partition, which reduces the overall search time.
+Popular tree-based methods include Vantage Point Tree (VP-tree) [25], M-tree[26] and Slim-Tree [27].
+More recently, several methods have been proposed to optimise the structure of the disk index, such as HD-Index [28], or methods that focus on parallel or distributed processing, such as Hercules [29] and TARDIS [30], respectively.
+Hashing-based methods aim to transform objects into low-dimensional representations, called hash codes [31].
+The hash codes are short sequences of bits designed to preserve the similarity between the original objects, ensuring that similar objects are mapped to similar codes.
+These methods are typically divided into data dependent and data independent approaches.
+Data dependent methods, based on machine learning or deep learning algorithms, use the underlying structure and distribution of data to design hash functions, resulting in more accurate and meaningful hash codes.
+Some methods based on this approach are BoostMap [32], Sequential Projection Learning Hashing (SPLH) [33], K-Means Hashing (KMH) [34] and Deep Hashing Network architecture for supervised hashing (DHN) [35].
+In contrast, data independent methods, such as LSH [4], use random projections to map samples in a feature space without requiring detailed knowledge of the data.
+Graph-based methods exploit the inherent structure of data by representing it as a graph, where nodes correspond to data points and edges represent the similarity relationship between the connecting pair of nodes [24].
+This approach facilitates the retrieval of similar objects through graph traversal algorithms, which can efficiently narrow down the search space.
+Several graph methods have been proposed, such as $k$-Nearest Neighbour Graph (KNNG) [36], NN-Descent [37], Hierarchical Navigable Small World Graph (HNSW) [38], Fast Approximate Nearest Neighbour Graph (FANNG) [39] and Navigating Spreading-out Graph (NSG) [40].
 2.3 Using different distance functions
@@
 The Manhattan or city block distance ($\ell_{1}$ norm), the Euclidean distance ($\ell_{2}$ norm), and the Chebyshev distance ($\ell_{\infty}$ norm) are typical instances of this family.
-As many multidimensional indexing structures were initially proposed in the context of spatial systems, many indexing structures and algorithms have commonly relied on the Euclidean distance, derived from its conventional use in two- or three-dimensional spatial applications.
-However, several negative consequences and challenges arise when dealing with high dimensional problems, an effect widely known as the Curse of Dimensionality.
+Many indexing structures and algorithms have commonly relied on the Euclidean distance, derived from its conventional use in two- or three-dimensional spatial applications.
+However, several drawbacks arise when dealing with high-dimensional problems, an effect widely known as the Curse of Dimensionality.
 As the number of dimensions increases, the volume of the space grows exponentially, and the density of points in the space decreases.
-This sparsity poses a challenge in accurately measuring distances and similarities between data points, since proximity becomes less meaningful.
-Consequently, points can be very far apart, even though they are close according to a metric such as Euclidean distance.
+This sparsity poses a challenge in accurately measuring distances and similarities between data elements, since proximity becomes less meaningful.
+Consequently, elements can be very far apart, even though they are close according to a metric such as the Euclidean distance.
 This effect increases the computational complexity of algorithms and operations performed in high-dimensional spaces, often resulting in slower and less efficient computations.
 For this reason, several authors have pointed out that, due to the Curse of Dimensionality, it is very costly to find the exact nearest neighbour in high-dimensional Euclidean spaces [43, 44].
-As a result, although Euclidean spaces reflect our natural conception of space, this natural assumption is challenging to extrapolate when dealing with high-dimensional data.
-Moreover, traditional clustering algorithms, such as $k$-means, often specify the Euclidean distance as the only appropriate metric without considering other possibilities.
+As a result, although Euclidean spaces reflect an intuitive conception of distance, it is challenging to extrapolate this natural assumption with high-dimensional data.
+Moreover, traditional clustering algorithms, such as $k$-means, often specify the Euclidean distance as the preferred metric without considering other possibilities.
 This choice brings in a self-imposed limitation that may produce results far from the best [45, 46, 9].
-In recent years, several surveys [23, 24, 25] have addressed the evaluation of various metrics for similarity searches across spaces of diverse topology, a burgeoning topic offering insights into the most suitable similarity measures for different problems and contributing to the ongoing development of new methods for similarity search.
-Therefore, we aim to introduce GMASK, an algorithm that, being based on a generalisation of earlier concepts previously introduced by MASK, signifies a substantial departure from the latter.
-GMASK enables the integration of arbitrary distances through the use of data partitioning algorithms that generate Voronoi regions and assign a prototype for each of these regions.
-Specifically, the utilisation of the $k$-medoids algorithm allows for the incorporation of multiple $\ell_{p}$ norms, facilitating parametrisation with different distance metrics and the execution of validation experiments.
+In recent years, several surveys [47, 48, 49] have addressed the evaluation of alternative metrics for similarity searches across spaces of diverse topology, a burgeoning topic offering insights into the most suitable similarity measures for different problems and contributing to the ongoing development of new methods for similarity search.
+Therefore, we aim to introduce GDASC, an algorithm that enables the integration of arbitrary distances through the use of data partitioning algorithms that generate Voronoi regions and assign a prototype for each of these regions.
+Specifically, the utilisation of several clustering algorithms, such as $k$-medoids, allows for the incorporation of multiple $\ell_{p}$ norms or even other metrics for specific contexts, facilitating parametrisation with different distances and the execution of validation experiments.
 Our findings demonstrate the feasibility of employing diverse distances beyond the Euclidean metric, highlighting the potential for achieving superior results in data indexing, contingent upon the specific characteristics of the dataset.
 3 Methodology
-This section describes the methodology developed for data indexation using arbitrary distance functions.
-First, it illustrates the GMASK algorithm as a general approximate indexing method accepting any distance function.
-Hereafter, it delves into the characteristics of Minkowski distances, emphasising the inherent limitations in $k$-means that justify the implementation of GMASK.
-Consequently, we justify adopting $k$-medoids, an alternative clustering method that leverages the opportunities provided by Minkowski distances to overcome previous restrictions derived from using the Euclidean distance.
-The design of experimental tests is described in Subsections “Performance evaluation”, “Benchmark algorithms for similarity search”, and “Datasets”, including metrics to assess performance, state-of-the-art algorithms for comparison, and the datasets selected to gauge effectiveness.
-3.1 GMASK
-As mentioned above, there is a close connection between machine learning and similarity search indexing.
-Firstly, the foundations for designing data access methods rely on familiar concepts in machine learning classification, such as dissimilarity spaces and distance functions.
-Secondly, clustering methods like $k$-means or $k$-medoids are used in many of these algorithms to create indexing structures for similarity search and improve search performance.
-Besides, they are also affected by the same limitations, such as the Curse of Dimensionality and the challenge of identifying the actual intrinsic dimensionality governing some problems to make them solvable.
-The typical choice in partitioning clustering methods is to find a low value of $k$ that generates different groups in the dataset.
-GMASK follows an unconventional approach for clustering, generating a high number of $k$ representative points for each level.
-The only requirement in GMASK is using a clustering method that generates Voronoi regions and returns a representative point for each of them.
-In this implementation, GMASK uses the $k$-medoids clustering method, inducing a Voronoi tessellation and returning an element representing each region.
-Thus, generating a large number of representative points, GMASK summarises the underlying data points with finer detail.
-In the same way, this indexing method applies to any dissimilarity space, including the particular case of metric spaces or vector spaces.
-Thanks to the $k$-medoids adaptation in GMASK, this algorithm presents an important improvement over MASK: it is able to work with different distance functions.
-This modification brings GMASK the capability of adaptation to the specific characteristics of any dataset.
-In the “Experimental results” section, we show that this unusual application of clustering has attractive properties and yields superior performance, even with high-dimensional datasets residing in feature spaces that are typically sparse.
-At this point, it is essential to note that the goal of GMASK is not to cluster data but to use the partitioning clustering method to construct a multilevel index to find elements.
-With more powerful computational infrastructures and large memory and storage capacity becoming prevalent, the advantage of this method lies in that previous restrictions on the number of clusters at each level become less relevant.
-In this new scenario, we can exploit the effect of introducing a high number of prototype points into the dataset, leveraging the capabilities of the computing infrastructure.
-Figure 2 illustrates the design principles of GMASK.
+This section describes a new data indexing framework called GDASC.
+First, it illustrates the properties and structure of GDASC.
+Then, it delves into the characteristics of different distance metrics, emphasising the inherent limitations in $k$-means.
+Aside from this traditional clustering algorithm, GDASC accepts other clustering algorithms such as $k$-medoids, which can exploit the properties of non-Euclidean or even non-metric distance functions for specific datasets.
+The design of experimental tests is described in Subsections 3.4, 3.5, and 3.6, including the metrics used to assess performance, state-of-the-art algorithms for comparison, and the datasets selected to gauge their effectiveness.
+3.1 GDASC
+GDASC is a versatile framework for data indexing as it can adopt different clustering algorithms and distance functions.
+Figure 2 depicts the three main blocks of the framework: Dataset Description (DD), Multilevel Structure Algorithm (MSA) and the Neighbours Search Algorithm (NSA).
+Besides the Dataset Description, GDASC relies on two algorithms, one for building a multilevel index structure (MSA) and another one for searching neighbours (NSA), described in Algorithm 1 and Algorithm 2 respectively.
+The proposed approach for building the multilevel structure, which is used in the search process, is prepared to receive any clustering algorithm that generates representatives summarising the underlying dataset.
+This is an important feature of GDASC since it is possible to adapt the framework to the particular characteristics of the dataset.
+Figure 2: GDASC framework is composed of three main parts.
+Given that one of the main contributions of GDASC is its applicability to real-world datasets with minimal or no preprocessing, the DD module performs an initial descriptive analysis of the dataset before conducting any experiments.
+Its goal is to determine adequate values for the input parameters of MSA and NSA.
+Firstly, by examining the number of elements and dimensions of the dataset, we can determine if the problem is described in a high-dimensional space or is computationally unfeasible.
+This assessment may lead us to consider distributing the dataset across multiple computational nodes.
+Next, for each dimension of the dataset, we will calculate the range of values to determine its sparsity and to gain an initial understanding of the distribution of points.
+We will also compute measures of central tendency (mean and median), dispersion (standard deviation), and distribution shape (kurtosis and skewness).
+Finally, for each distance metric to be applied for similarity search, we calculate the maximum and minimum distances between two points in the dataset, the average distance between points, and the distribution of distance quantiles.
+These calculations provide a deeper understanding of the dataset and help determine the values of initial parameters that depend on dataset features, such as the radius.
+The MSA module uses Algorithm 1 to build a multilevel index structure following a bottom-up approach and using a cluster algorithm with an initial number of prototypes as high as possible, depending on the computational capacity.
+In the use of clustering algorithms, is very common to find a low number of centroids that generates different groups in the dataset.
+However, MSA follows an rather unusual approach for clustering, generating instead a high number of representative points for each level (prototypes).
+This way, the resulting indexing structure can summarise the underlying data points with finer detail.
+The structure generated by the MSA module depends on several initial parameters:
+  • The group length ($gl$) quantifies the number of elements in each data partition at the lowest level, configured according to the available computing power (and the total number of nodes in distributed computing systems).
+  • The number of prototypes ($nPrototypes$) calculated for each group.
+Contrary to the standard procedure, our approach suggests that, for information indexing, one should use as many prototypes as the computational infrastructure can afford.
+Empirically we have concluded that the ratio between the group length and the number of prototypes have to be 2:1.
+This ratio also determines the number of layers of the multilevel index.
+  • The distance function ($distFunc$) applied by the clustering algorithm. This is a critical decision because not all distance functions work properly with any dataset.
+  • The clustering algorithm ($clusterAlg$).
+Clustering methods like $k$-means or $k$-medoids are widely used in algorithms that create indexing structures for similarity search.
+However, MSA is able to work with any clustering algorithm.
+In this paper, we focus in the use of $k$-medoids and consequently employ various distance functions.
+Algorithm 1 Multilevel Structure Algorithm in GDASC
+1:procedure MSA($data,gl,nPrototypes,distFunc,clusterAlg$)
+2:     $vector,nGroups\leftarrow$ split($data,gl$); $\triangleright$ Divide dataset into groups of $gl$
+3:     $\triangleright$ Initialise multilevel structure: $levelPoints$, $levelLabels$, and variable $idLevel$
+4:     while $nGroups\geq 1$ do
+5:         $\triangleright$ Initialise auxiliary structures for each level: $groupsPoints$, $groupsLabels$ and $points$
+6:         for $idGroup=1,nGroups$ do
+7:              $points\leftarrow vector[idGroup]$;
+8:              $groupsPoints.add($clustering($clusterAlg,distFunc,nPrototypes,points$));
+9:              $groupsLabels.add(clusters.labels)$;
+10:         end for
+11:         $levelPoints.add(groupsPoints)$;
+12:         $levelLabels.add(groupsLabels)$;
+13:         $vector,nGroups\leftarrow$ split($levelPoints,gl$);
+14:         $idLevel\mathrel{+}=1$;
+15:     end while
+16:     $nLevels\leftarrow idLevel-1$;
+17:end procedure
+Figure 3 illustrates the schema implemented by MSA.
 At the lowest level, we have the data points in $\mathcal{X}$ to be indexed.
-The blue boxes in the diagram represent different data partitions or groups that could be stored on various computing nodes.
-The novelty lies in creating the multilevel structure of $k$ medoids, following a bottom-up approach.
-The first level of medoids (identified as $n_{1}$ in Figure 2) summarises the data points at the lowest level so that points assigned to each cluster are represented by its prototype (medoid).
-The second level (denoted $n_{2}$) corresponds to medoids representing the prototypes computed at the first level, $n_{1}$.
-The same strategy is repeated, creating additional prototype layers that summarise the immediately lower group of medoids.
-The procedure stops when the number of medoids at the top level of this index structure becomes manageable.
-Figure 2: Conceptual representation of the updated multilevel structure approach created by GMASK.
-The process of constructing this multilevel index structure depends on two initial parameters:
-  • The group length quantifies the number of elements in each data partition at the lowest level, configured according to the available computing power (and the total number of nodes in distributed computing systems).
-For the baseline, we assume that all partitions are of the same size, although this condition is not strictly necessary.
-  • The number of medoids calculated for each group.
-At this point, the question of selecting the most optimal number of medoids has not yet been addressed, but it is not very important.
-Contrary to the standard procedure, our approach suggests that, for information indexing, one should use as many medoids as the computational infrastructure can afford.
-The clustering at the first level reduces the initial number of data points by a particular proportion, the data summarisation ratio, which is determined by the ratio between the group length and the number of medoids.
-This ratio also determines the number of layers of the multilevel index and the stop condition when the number of medoids at the top layer becomes lower than the group length.
-Each medoid is a prototypical point representing all the elements assigned to its cluster.
-The number of points assigned to each cluster may vary depending on their distance from the corresponding medoid.
-For this reason, regions with higher data density will also tend to receive more medoids.
-By generating a high number of medoids, we can capture finer details of the underlying data, allowing us to index not only high-density regions but also outliers and clusters of elements distinct from the main concentrations of points.
-Algorithm 1 describes the construction of this multilevel index structure.
-Algorithm 1 Multilevel index construction in GMASK
-1:function Clustering($distFunc,nPrototypes,points$)
-2:     $\triangleright$ A clustering method returning Voronoi regions
-3:     $\triangleright$ and a prototype point representing each region
-4:end function
-5:
-6:procedure GMASK($data,lengthGroup,nPrototypes$)
-7:     $ngroups\leftarrow lengthData/lengthGroup$;
-8:     $vector\leftarrow$ split($data,nGroups$); $\triangleright$ Divide dataset into nGroups
-9:     $\triangleright$ Initialise lists $layerPoints$, $layerLabels$, and variable $idLayer$
-10:     while $ngroups\geq 1$ do
-11:         $\triangleright$ Initialise groupsPoints, groupsLabels and points
-12:         for $idGroup=1,nGroups$ do
-13:              $points\leftarrow vector[idGroup]$;
-14:              $groupsPoints.add($clustering($distFunc,nPrototypes,points$));
-15:              $groupsLabels.add(clusters.labels)$;
-16:         end for
-17:         $layerPoints.add(groupsPoints)$;
-18:         $layerLabels.add(groupsLabels)$;
-19:         $ngroups\leftarrow length(groupsPoints)/lengthGroup$;
-20:         if $ngroups\geq 1$ then
-21:              $vector\leftarrow$ split($layerPoints,nGroups$);
-22:         end if
-23:         $idLayer\mathrel{+}=1$;
-24:     end while
-25:     $layers\leftarrow idLayer-1$;
-26:end procedure
-Once built according to the rules described above, the index is ready to accept point queries, which systematically search for all points in the original dataset, seeking a perfect match.
-To solve any query using this multilevel index, we start at the top level of the medoids hierarchy, steering the search according to the distance between the medoids at each level and the target query object $q\in\mathcal{X}$.
-3.2 Dissimilarity measures
-A dissimilarity space [26] defines the rules for measuring the distance between elements represented in a feature space.
-As discussed in Section “Similarity search”, similarity captures the notion of closeness between two objects.
-Euclidean vector spaces are widely applied in previous research, as they reflect and intuitive interpretation of a space: vectors are elements that can be added, multiplied by a scalar or projected onto each other.
-Being $\mathbb{X}$ a set of elements, the function
-$\delta:\mathbb{X}\times\mathbb{X}\to\mathbb{R}$ (1)
-returns a real number $\delta(a,b)$ that shall be used to quantify the difference between two or more elements contained on this set, denoting a measure of distance.
-Distance functions have the following properties:
-  • $(p1)\quad\forall a,b\in\mathbb{X},\delta(a,b)\geq 0$ positiveness,
-  • $(p2)\quad\forall a,b\in\mathbb{X},\delta(a,b)=\delta(b,a)$ symmetry,
-  • $(p3)\quad\forall a\in\mathbb{X},\delta(a,a)=0$ reflexivity,
-  • and in most cases
-  • $(p4)\quad\forall a,b\in\mathbb{X},a\neq b\to\delta(a,b)>0$ strict positiveness.
-If it also satisfies
-  • $(p5)\quad\forall a,b,c\in\mathbb{X},\delta(a,b)\leq\delta(a,c)+\delta(c,b)$, commonly known as triangle inequality,
-$\delta$ is indeed a metric, and the pair $(\mathbb{X},\delta)$ is called a metric space [6].
-If the distance does not satisfy the strict positiveness property $(p4)$ then the space is called a pseudo-metric space, as well as if it does not hold the symmetry property $(p2)$, we may have a quasi-metric space.
-Nevertheless, there are some primitive spaces where these metric constraints are not necessarily satisfied.
-We also use the word ’dissimilarity’ when referring to this difference or distance between elements [18].
-Normed and inner product spaces are special cases of metric spaces, where a metric is defined either by a norm or an inner product.
-An Euclidean space is a specific case of a Hilbert space.
-In turn, a Hilbert space is an inner product space and a particular case of a Banach space, an example of normed spaces in which a norm metric $\ell norm$ has been defined [5].
-3.3 Minkowski distance
-Minkowski distances or $\ell_{p}norms$ are distance measures commonly used to compute the dissimilarity of objects described by numerical attributes.
-However, dissimilarity can also be quantified using non-numerical attributes through other alternatives, such as the edit distance widely used for text attributes.
+They are divided into several groups according to the $gl$ parameter.
+Once these groups are formed, the points in the next level are the representative points of the previous level.
+That is, following a bottom-up approach, the first level of prototypes (identified as $n_{1}$ $prototypes$ in Figure 3) summarises the data points at the lowest level so that points assigned to each cluster are represented by its prototype.
+The second level (denoted $n_{2}$ $prototypes$) corresponds to prototypes representing the prototypes computed at the first level, $n_{1}$ $prototypes$.
+The same strategy is repeated, creating additional prototype layers that summarise the immediately lower group of prototypes.
+The procedure stops when the number of prototypes at the top level of this index structure becomes manageable.
+It can be remarkable that the groups formed at the lower level or the cluster formed at any level can have distinct size.
+When the number of points in a group or a cluster is less or equal than the number of prototypes, all the points are become prototypes.
+In the next level, the groups or prototypes are gather in groups with, at least, as points as $gl$.
+At level 0, the last three groups have less points than $gl$.
+Thus, in level 1, the prototypes of these groups are arranged in one group and five prototypes are selected for level 2.
+Figure 3: Conceptual representation of the updated multilevel structure approach created by GDASC.
+The points assigned to each cluster may vary depending on the distance function applied and its value.
+For this reason, regions with higher data density will also tend to receive more prototypes.
+By generating a high number of prototypes, we can capture finer details of the underlying data, allowing us to index not only high-density regions but also outliers and clusters of elements distinct from the main concentrations of points.
+The NSA module uses the multilevel index structure generated by MSA to find the $k$ nearest neighbours by applying Algorithm 2.
+For a given point query $(q)$, a number of neighbours $(k)$ and a radius, this algorithm goes from the top level of the multilevel structure to the bottom one, testing the distance between prototypes and the query point.
+In each level, the algorithm computes the distance between the point query and the prototypes in that level.
+Only the prototypes whose distance is less than radius will be consider in the next level.
+Therefore, the search way goes throw prototypes that satisfy this filtering criterion.
+Algorithm 2 Neighbours Search Algorithm in GDASC
+procedure NSA($levelPoints,levelLabels,nLevels,distFunc,q,k,radius$)
+2:     $prototypes\leftarrow levelPoints[nLevels]$;
+     $D\leftarrow$ distance($distFunc,q,prototypes$);
+4:     $idCandidates\leftarrow D<radius$;
+     for $level=nLevels-1,1$ do
+6:         $prototypes\leftarrow levelPoints[levels][idCandidates]$;
+         $D\leftarrow$ distance($distFunc,q,prototypes$);
+8:         $idCandidates\leftarrow D<radius$;
+         $level\mathrel{1}=1$;
+10:     end for
+     $candidates\leftarrow levelPoints[0][idCandidates]$;
+12:     $D\leftarrow$ distance($distFunc,q,candidates$);
+     $candidates\leftarrow$ sort($candidates,D$);
+14:     $neighbours\leftarrow candidates[0:k]$;
+end procedure
+Figure 4 represents the search procedure of $k=2$ neighbours using a multilevel structure with 4 levels.
+In a previous step, NSA computes the distance between the point query $q$ and the prototypes at level 3, $d(C_{3-1},q),d(C_{3-2},q),d(C_{3-3},q)$.
+As only $d(C_{3-1},q)$ and $d(C_{3-3},q)$ are less than radius, only the points in $C_{3-1}$ and $C_{3-2}$ clusters are considered.
+Thus, the algorithm calculates the distance between $q$ and all the points in $C_{3-1}$ and $C_{3-3}$ clusters and only selects the points for which this value is less than radius.
+Paying attention to $C_{3-1}$, this cluster contains three points and only the distance for two of them are less than radius.
+Therefore, the arrows between $C_{3-1}$ and these points are represented by a solid pink line, while the other arrow is represented by a dotted grey line.
+In the next level, the algorithm only analyses the distances for the points in $C_{2-1}$ and $C_{2-3}$ clusters.
+When the bottom level is reached, the algorithm has selected several points as potential candidates.
+They are filtered according to the distance between them and $q$ selecting the $k=2$ points with less distance value.
+Figure 4: Search algorithm example.
+3.2 Distance metrics
+Distance metrics provide a quantitative means of assessing the dissimilarity between objects represented in a feature space.
+In this paper we focus on two specific distance metrics: Minkowski distance, which generalises multiple distance measures, and cosine distance, which assesses the dissimilarity between vectors based on their angle in a high-dimensional space.
+Minkowski distance
+The Minkowski distance or $\ell_{p}norm$ conforms a set of distance measures commonly used to compute the dissimilarity of objects described by numerical attributes.
+However, dissimilarity can also be quantified using non-numerical attributes through other alternatives, such as the edit distance, widely used for text attributes.
 Let $\ell_{\infty}^{p},p\geq 1$ be a vector space of real sequences such that $\sum_{i=1}^{\infty}|x_{i}|^{p}$ with the norm given by $||x||_{p}(\sum_{i=1}^{\infty}|x_{i}|^{p})^{\frac{1}{p}}$.
@@
 Nonetheless, several authors have shown that fractional distances can significantly improve the effectiveness of standard clustering algorithms such as $k$-means [46].
-3.4 Limitations of k-means and the Euclidean distance
-In unsupervised machine learning with $k$-means, the user must determine the appropriate number of clusters for a given problem, and a centroid represents each cluster.
-Carefully choosing the number of centroids is crucial because too many centroids would not effectively summarise the underlying data, and too few centroids would result in grouping unrelated data points.
-However, the locations provided to these initial centroids can greatly impact the final outcome, making it a sensitive aspect.
-To address this limitation, algorithms such as $k$-means++ [27] propose an approach where the initial random centroids are spread more evenly, leading to improved results.
-Nevertheless, although several more powerful $k$-means implementations have tried to reduce the performance impact resulting from sub-optimal initial centroid placement, only a few have explored the potential for achieving better solutions by choosing a more appropriate distance.
-Due to the nature of the $k$-means algorithm, the choice of distance function can also exert a significant impact on $k$-means outcomes, as it determines how similarity between data points is calculated and produces clusters with different shapes and separations between each other.
-As it has been raised in Section “Background and related work”, despite the widespread use of Euclidean distance in indexing structures and algorithms, relying on it as the default option when using the $k$-means algorithm can lead to poor results.
-Previous research [22] suggests that inherent characteristics of high-dimensional spaces, such as high dispersion, outliers or concentration of points around specific values, hinder the discriminatory power of the $\ell_{p}norm$ for higher values of $p$.
-This situation would provide undue importance to the distant (sparse/noisy) dimensions and poorly discriminate between further and nearer neighbours.
-Therefore, for a given high-dimensional problem, it may be preferable to use Minkowski distances with lower values of $p$, so the $\ell_{1}$ distance metric (Manhattan distance) would be the best choice for high-dimensional applications, followed by the Euclidean distance ($\ell_{2}$), then the $\ell_{3}$ metric, up to the Chebyshev distance ($\ell_{\infty}$).
-Previous algorithms for approximate similarity search, such as MASK [2], have relied on the Euclidean distance to calculate relative dissimilarities between points.
+Cosine distance
+The cosine distance is a measure of dissimilarity between two non-zero vectors in an inner product space, which calculates the cosine of the angle between them.
+It is a handy distance in high-dimensional spaces and is widely applied in many fields such as text mining, information retrieval, and machine learning, especially in document similarity and clustering [23].
+Unlike the Euclidean distance that considers the magnitude of the vectors, cosine distance focuses on their orientation, making it less sensitive to differences in vector length and more robust to variations in data scaling [46].
+Cosine distance has two main differences from Euclidean distance.
+Firstly, it is scale-invariant because it measures the angle between two vectors rather than their magnitude.
+Secondly, it assigns higher weight to dimensions with higher values, whereas Euclidean distance treats all dimensions equally.
+The cosine distance measures the dissimilarity between two non-zero vectors and can be formulated as one minus the cosine similarity between them.
+This property makes it particularly effective in handling sparse data and datasets with many features, where traditional distance metrics often struggle due to the Curse of Dimensionality.
+Mathematically, for two vectors A and B, the cosine distance is defined as:
+$Distance_{A,B}=1-cos(\theta)=1-\frac{A\cdot B}{\|A\|\|B\|}=1-\frac{\sum_{i=1}^{n}A_{i}B_{i}}{\sqrt{\sum_{i=1}^{n}A_{i}^{2}}\sqrt{\sum_{i=1}^{n}B_{i}^{2}}},$ (2)
+where $\theta$ is the angle between A and B in the n-dimensional space, and $A_{i}$ and $B_{i}$ are the $i$-th components of vectors A and B, respectively.
+Thus, the cosine distance ranges from 0 to 2, where 0 means that the vectors are identical and 2 means that they are diametrically opposed to each other.
+3.3 Clustering algorithms
+There are many clustering algorithms for unsupervised machine learning and one of the most common options is $k$-means.
+When applying $k$-means the user must usually determine the appropriate number of clusters for a given problem, and the algorithm obtains a centroid representing each cluster.
+Choosing the number of clusters is crucial because too many centroids would not effectively summarise the underlying data and too few centroids would result in grouping unrelated data points.
+Besides, the locations assigned to these initial centroids can greatly impact the final outcome, making it a sensitive aspect.
+To elude this limitation, algorithms such as $k$-means++ [50] propose an approach where the initial random centroids are spread more evenly, leading to improved results.
+Nevertheless, although several more powerful $k$-means implementations have tried to reduce the performance impact resulting from sub-optimal centroid initialisation, only a few have explored the potential for achieving better solutions by choosing a more appropriate distance.
+The choice of distance function to be used with $k$-means, can also exert a significant impact since it determines how similarity between data points is calculated and produces clusters with different shapes and separations.
+Previous research [46] suggests that inherent characteristics of high-dimensional spaces, such as high dispersion, outliers or high concentration of points around specific values, hinder the discriminatory power of the $\ell_{p}norm$ for higher values of $p$.
+This situation would provide undue importance to the distant (sparse/noisy) dimensions and poorly discriminate between further and nearer neighbours, making it preferable to use Minkowski distances with lower values of $p$.
+For instance, the $\ell_{1}$ distance metric (Manhattan distance) would be a better choice for high-dimensional applications, followed by the Euclidean distance ($\ell_{2}$), then the $\ell_{3}$ metric, and so for, up to the Chebyshev distance ($\ell_{\infty}$).
+As we have indicated in Section 2, despite the widespread use of indexing structures and algorithms based on the Euclidean distance [51], as a default option when using the $k$-means algorithm, such a choice can lead to poor results.
 This strategy follows the conventional inclination to associate distance between points with the Euclidean metric, which aligns with the distance concept commonly applied in geospatial problems.
-However, the potential for achieving improved results, coupled with the lack of prior empirical studies that validate this proposal and the need for learning paradigms applicable to a broader range of dissimilarity measures [5], justify our efforts to explore alternative Minkowski distances for data indexing.
-This endeavour aims to foster a deeper understanding of the impact of using alternative distance metrics with high-dimensional data.
+However, the potential for achieving improved results, coupled with the lack of prior empirical studies validating this proposal, and the need for learning paradigms applicable to a broader range of dissimilarity measures [9], justify our efforts to explore non-Euclidean distances for data indexing.
+This endeavour fosters a deeper understanding of the impact of using alternative distance functions with high-dimensional data.
 Nevertheless, while it may seem theoretically plausible to employ $k$-means or its enhanced version, $k$-means++, with several distance metrics, the practical implementation reveals that the algorithm’s internal design is intertwined with the Euclidean distance.
-This alignment relies on the algorithm’s iterative optimisation process when it searches for centroids that minimise the total squared distances of points within their respective clusters[3].
-The sum of the Squared Error (SSE), calculated as the sum of squared Euclidean distances between each data point and the centroid of its assigned cluster, underscores the algorithm’s inherent reliance on Euclidean geometry for clustering optimisation [28].
-As we have already demonstrated, assuming that widely adopted mechanisms are the most suitable or the only viable option can lead to self-imposed limitations that generate sub-optimal outcomes.
-Instead, our objective is to steer clear of the misconception that certain implementations are the best or the only conceivable choice simply because they are the most common choices.
-At this juncture, GMASK can integrate an alternative clustering algorithm that leverages the opportunities provided by diverse distance metrics for data indexing.
-The $k$-means algorithm defines clusters based on centroids minimising the sum of squared errors in distances from all other points in the same cluster.
-In contrast, alternative methods like $k$-medoids calculate clusters around the most representative data element, which is determined by minimising the sum of pairwise dissimilarities among data points the same cluster.
-3.5 k-medoids as a generalisation of k-means
-Several generalisations of the $k$-means algorithm have been proposed to broaden the scope of applicable data types [29], reflecting the diversity found in nature prototype points, such as subsets of objects, hyperplanes, etc.
-The first paper proposes a general criterion and its generalised $k$-means method is [30].
+This alignment relies on the algorithm’s iterative optimisation process, when searching for centroids that minimise the total squared distances of points within their respective clusters[52].
+The sum of the Squared Error (SSE), calculated as the sum of squared Euclidean distances between each data point and the centroid of its assigned cluster (so-called Within-Cluster Sum of Squares or WCSS), underscores the algorithm’s inherent reliance on Euclidean geometry for clustering optimisation [53].
+As we have already demonstrated, assuming that widely adopted mechanisms are the most suitable or the only viable option can lead to sub-optimal outcomes.
+Instead, our goal is to steer clear of the misconception that certain implementations are the best or the only conceivable choice because they are the most common.
+Thus, GDASC can integrate alternative clustering algorithms to create the indexing structure.
+While $k$-means defines clusters based on centroids minimising the sum of squared errors in distances from all other points in the same cluster, alternative methods like $k$-medoids calculate clusters around the most representative data element, which is determined by minimising the sum of pairwise dissimilarities among data points in the same cluster.
+The first paper proposing a general criterion and a generalised $k$-means method is [54].
 This approach starts from a $n$ x $n$ dissimilarity matrix $(d_{\ell}t)$ for $n$ factories $O=1,...,n$ in an industrial network where $d_{\ell}t$ are the minimum road transportation costs between $\ell$ and $t$.
@@
 $Q(C_{i},z):=\sum_{\ell\in C_{i}}d(\ell,z)\to\min_{z\in C_{i}}$ (3)
-Many authors, including, among others, Diday [31, 32, 33], Bock [34] or Govaert [35], have followed the generalised clustering approach in various settings and numerous variations, obtaining a plethora of generalised $k$-means algorithms.
-Kaufmann and Rousseeuw [36] first propose the popular Partitioning Around Medoids (PAM) algorithm, also referred to as $k$-medoids.
-In $k$-medoids, data is modelled by using $k$ representative objects $m_{i}$ called medoids, that serve as prototypes for the clusters (formerly called classes) which are obtained by using the absolute error criterion (’total deviation’, TD).
-Since the TD is defined as the sum of dissimilarities of each point $x_{j}\in C_{i}$ to the medoid $m_{i}$ of its cluster, the medoid of a cluster $C_{i}$ is in fact the object with the smallest sum of dissimilarities (or equivalently, the smallest average) to all other objects in the cluster.
+Many authors, including Diday [1971, 1973, 1974], Bock [1970] or Govaert [59], have followed the generalised clustering approach in various settings and numerous variations, obtaining a plethora of generalised $k$-means algorithms.
+Kaufmann and Rousseeuw [1987] first proposed the popular Partitioning Around Medoids (PAM) algorithm, also referred to as $k$-medoids.
+Data is modelled using $k$ representative objects $m_{i}$ (called medoids), that serve as prototypes for the clusters (formerly called classes).
+These prototypes are obtained using the absolute error criterion (’total deviation’, TD).
+Since the TD is defined as the sum of dissimilarities of each point $x_{j}\in C_{i}$ to the medoid $m_{i}$ of its cluster, the medoid of a cluster $C_{i}$ is in fact the object with the smallest sum of dissimilarities.
 $medoid(C_{i}):=argmin_{x_{i}\in C}\sum_{{x_{j}\in C}}d(x_{i},x_{j})$ (4)
 This method is suitable for non-Euclidean distances and a wide range of data since it does not force the dissimilarity measure to adhere strictly to metric properties, and only requires a proximity measure between any pair of objects.
-In addition, while $k$-means defines a prototype in terms of a centroid, which is usually the mean of a group of points that may not correspond to an element in the dataset, $k$-medoids defines a prototype in terms of a medoid, an actual data point identified as the most representative one for a group of points.
-Regarding $k$-medoid implementations, Program PAM [36] consists of two algorithms, BUILD to choose an initial clustering, and SWAP to improve the clustering towards a local optimum, since finding the global optimum of the $k$-medoids problem is, unfortunately, NP-hard.
-The algorithms require a dissimilarity matrix, which requires $O(n^{2})$ memory and at least $O(n^{2}d)$ time to be computed.
+While $k$-means defines a prototype in terms of a centroid, $k$-medoids defines a prototype in terms of a medoid, an actual data point identified as the most representative one for a group of points.
+As for $k$-medoids implementations, Program PAM [60] consists of two algorithms: BUILD to choose an initial clustering and SWAP to improve the clustering towards a local optimum, since finding the global optimum of the $k$-medoids problem is NP-hard.
+The algorithms need a dissimilarity matrix, which requires $O(n^{2})$ memory and at least $O(n^{2}d)$ time to be computed.
 On top of this, the process of updating medoids involves evaluating the cost of swapping non-medoid points to ensure that the sum of pairwise dissimilarities within clusters (rather than the sum of squared Euclidean distances) is minimised, making this method more robust to noise and outliers compared to the alternatives based on $k$-means.
 Recently, FasterPAM [61] introduced a speedup for larger $k$, achieving an $O(k)$-fold acceleration in the second SWAP phase of the algorithm, but still providing the same results as the original PAM algorithm.
-Originally FasterPAM was implemented in Java and published within the open-source library ELKI [38].
-However, to integrate it with GMASK implementation, developed entirely in Python, we will employ the fast $k$-medoids Python implementation [39] version provided by the $kmedoids$ package  [^1] https://pypi.org/project/kmedoids.
+Originally, FasterPAM was implemented in Java and published within the open-source library ELKI [62].
+However, to integrate it with GDASC implementation, developed entirely in Python, we will employ the fast $k$-medoids Python implementation [63] version provided by the $kmedoids$ package  [^1] https://pypi.org/project/kmedoids.
 FasterPAM let us upgrade our method by making it compatible with any distance function and a wider range of problems.
-Experimental results verify the applicability of GMASK with real datasets, improving the performance of alternative algorithms and providing better results for approximate similarity search queries.
 3.4 Performance evaluation
-The main advantage of GMASK is providing a fast response to queries at the expense of returning approximate results.
-Thus, a straightforward strategy to assess the performance of this algorithm is to undertake an exhaustive search of all elements in a given dataset $X$.
-Then, the proportion of data points that were correctly retrieved can be determined.
-This is the common approach to all evaluation experiments developed in this study.
-A comprehensive series of experiments are conducted across datasets of diverse types, sizes, dimensionality, and sparsity to evaluate the performance of GMASK, which introduces $k$-medoids for the bottom-top construction of the index tree.
-This novelty enables its parametrisation with several $\ell_{p}$ norms as outlined in “k-medoids as a generalisation of k-means”.
-These results will be compared with those obtained using other state-of-the-art algorithms.
-As mentioned above, our objective is to understand better the impact of utilising distances other than the Euclidean metric.
-Through these experiments, we aim to demonstrate that data indexing using non-Euclidean distances is not only feasible but also has the potential to yield improved outcomes.
-3.7 Benchmark algorithms for similarity search
-To evaluate our thesis, we compare GMASK performance in approximate nearest neighbour search with respect to alternative MAM.
-3.7.1 KD-Tree
-The KD-Tree (K-Dimensional Tree) [40] is an exact similarity search method used to find the nearest neighbours to a given query point.
+A straightforward strategy to evaluate the performance of an algorithm is to search for the $k$ approximate nearest neighbours of every query point in a test dataset against a previously indexed train dataset.
+To measure the accuracy of the algorithm’s performance in this task and thereby assess its ability to retrieve the relevant elements during a search, we would use the recall metric.
+This quality measure is defined as:
+$Recall=\frac{\textit{Number of Relevant Items Retrieved}}{\textit{Total Numberof Relevant Items in the Search}}$
+In this context, the Relevant Items are defined as the ground truth neighbours, corresponding to the true or correct set of $k$ points that the algorithm should ideally identify as the nearest neighbours of a given query point.
+They are obtained through the use of an exact search algorithm that will be chosen based to the distance metric employed, depending on its compatibility with the algorithm:
+  • Brute-Force represents the most naive exact neighbour search implementation involves the computation of distances between all pairs of points in the dataset using a chosen metric (e.g.
+Euclidean distance, cosine similarity) and selecting the closest ones.
+Efficient brute-force neighbours searches can be very competitive for small data samples.
+However, as the number of samples grows, this approach quickly becomes infeasible.
+  • The KD-Tree (K-Dimensional Tree) [64] serves as a widely used method for exact similarity search.
 It constructs a binary tree structure by recursively dividing the data based on the maximum spread along different dimensions and the search algorithm efficiently narrows down the search space by traversing the tree.
@@
 However, as the dimensionality increases, the performance of the KD-tree can degrade due to the Curse of Dimensionality phenomenon.
-As it performs an exact search, the solutions provided by this method are used as a baseline to evaluate the performance of other approximate algorithms.
-We shall compare the k-nearest neighbours obtained by each method against those obtained using the KD-Tree to assess their accuracy.
-3.7.2 FLANN
-FLANN (Fast Library for Approximate Nearest Neighbours) [11] is an efficient method widely employed for approximate similarity search in large-scale and high-dimensional datasets, extensively used in machine learning, computer vision, and data mining.
+Although it is the preferred algorithm for calculating ground truth neighbours due to its higher computational efficiency, it supports a smaller range of distance metrics compared to the Brute Force method.
+3.5 Benchmark algorithms for similarity search
+After defining our evaluation strategy, we shall compare the performance of GDASC in approximate nearest neighbour search against alternative MAM to validate our thesis:
+  • FLANN (Fast Library for Approximate Nearest Neighbours) [15] is an efficient method widely employed for approximate similarity search in large-scale and high-dimensional datasets.
+This library is extensively used in machine learning, computer vision and data mining.
 FLANN employs techniques such as random projection and hierarchical subdivision to construct an index structure that enables faster search operations.
-It offers several index types, including KD-tree, $k$-means tree, and hierarchical clustering tree, which are suitable for diverse data types and query requirements.
+It offers several index types, including KD-tree, $k$-means tree and hierarchical clustering tree, which are suitable for diverse data types and query requirements.
 While FLANN has implementations in several programming languages, we use the Python bindings provided by the pyflann  [^2] https://github.com/primetang/pyflann library.
-3.7.3 PyNNDescent
-The PyNNDescent algorithm is an approximate similarity search method [41] implemented in Python  [^3] https://pynndescent.readthedocs.io/en/latest/.
-It is based on the NNDescent algorithm and used for efficient nearest neighbour search in large-scale datasets.
-PyNNDescent constructs a graph-based index by connecting each data point to its approximate nearest neighbours.
-It iteratively improves the graph structure through neighbour descent steps, employing techniques such as random initialisation, multi-tree searching, and pruning.
+  • The PyNNDescent (PyNN) algorithm is an approximate similarity search method [37] implemented in Python  [^3] https://pynndescent.readthedocs.io/en/latest/.
+It is based on the NNDescent algorithm and is used for efficient nearest neighbour search in large-scale datasets.
+PyNN constructs a graph-based index by connecting each data point to its approximate nearest neighbours.
+It iteratively improves the graph structure through neighbour descent steps, employing techniques such as random initialisation, multi-tree searching and pruning.
 The algorithm also provides parameters to control the trade-off between search accuracy and computational cost.
-PyNNDescent is particularly useful for high-dimensional datasets and is widely used in machine learning, data mining, and information retrieval tasks.
+PyNN is particularly useful for high-dimensional datasets and is widely used in machine learning, data mining and information retrieval tasks.
 3.6 Datasets
-Several experiments have been conducted to evaluate the performance of GMASK, using four datasets diverse in nature and public access. Table 1 briefly summarises their main features.
-3.8.1 Municipalities
+Several experiments have been conducted to evaluate the performance of GDASC using four publicly accessible datasets that vary in type, size, dimensionality (Dim), and sparsity (HS), with minimal or no preprocessing.Table 1 briefly summarises their main features.
+Dataset Label n Dim HS Data Type Purpose
+Municipalities $\mathcal{X}_{muni}$ 8,130 2 No Geospatial Outliers
+MNIST $\mathcal{X}_{MNIST}$ 69,000 784 Yes Image High Dimensionality & Dispersion
+GLOVE $\mathcal{X}_{GLOVE}$ 1,000,000 100 No Text High Number of Elements
+NYtimes $\mathcal{X}_{NYtimes}$ 290,000 256 No Text Cosine Distance
+Table 1: Descriptive traits characterising the datasets included in experiments.
+Municipalities
 The Municipalities dataset is a two-dimensional geospatial dataset that contains the coordinates (latitude and longitude) of each of the 8,130 municipalities that make up the Spanish territory, as depicted in Figure 5.
@@
 Figure 5: Visualisation for dataset Municipalities
-Although this is a two-dimensional dataset with a low dispersion, it is an exceptionally intriguing case due to its intuitive representation and the presence of outliers, such as the Canary Islands and the Balearic Islands.
-The location of these islands deviates from the rest of the points significantly and poses a challenge to any similarity search method applied to this dataset.
-3.8.2 MNIST
+In Figure 5, we can identify three main groups of points: the largest region corresponds to the peninsular territory of Spain, whereas the group of points located at the right of the map corresponds to the Balearic Islands and the group of points located at the bottom left corner of the map corresponds to the Canary Islands.
+As the location of these islands deviates from the rest of the points as significantly, we can consider them as outliers.
+Thus, despite being a two-dimensional dataset with a low dispersion, it is an exceptionally interesting case due to its intuitive representation and the presence of outliers, posing a significant challenge to any similarity search method applied.
+MNIST
 The MNIST_784 dataset  [^5] http://yann.lecun.com/exdb/mnist/ is a widely used machine learning benchmark dataset whose simplicity makes it an ideal starting point for algorithm development and evaluation.
@@
 Moreover, each image is 28x28 pixels in size, resulting in a total of 784 dimensions when flattened, allowing researchers to explore algorithms in the context of high-dimensional data.
-Many well-established techniques and algorithms have been evaluated on MNIST [42, 43], making it a reference point for evaluating the performance of approximate similarity search methods [16, 44].
-3.8.3 GLOVE
+Many well-established techniques and algorithms have been evaluated on MNIST [65], making it a reference point for evaluating the performance of approximate similarity search methods [66, 49].
+GLOVE
 The third experiment involves the GLOVE_100 dataset [^6] https://nlp.stanford.edu/data/glove.6B.zip, which contains one million pre-trained Global Vectors for Word Representation (GloVe) embeddings that capture semantic relationships between words.
@@
 Given their widespread adoption as a benchmark in the natural language processing community, the use of GloVe embeddings as a standardised dataset provides valuable insights into the method effectiveness in retrieving similar items based on their semantic similarity [66].
-3.8.4 NYtimes
+NYtimes
 The NYtimes dataset, part of the Bags of Words dataset from the UCI repository [^7] https://archive.ics.uci.edu/dataset/164/bag+of+words, comprises a collection of New York Times news articles represented as a bag of words.
@@
 The dataset is organised as a document–word matrix, where each row corresponds to a document and each column corresponds to a word.
-The values in the matrix indicate the frequency of each word occurring in the respective document.
+The values in the matrix indicate the frequency of each word occurs in the respective document.
 Preprocessing steps include tokenization, removal of stopwords, and vocabulary truncation, with only words occurring more than ten times retained.
-The reduced version of the NYtimes employed on this work was generated by building tf-idf descriptors from the bag-of-words version, and embedding them into a lower dimensional space using the Johnson-Lindenstrauss Transform [45].
-Dataset Label n Dimensionality High Sparsity Data Type
-Municipalities $\mathcal{X}_{muni}$ 8,130 2 No Geospatial
-MNIST $\mathcal{X}_{MNIST}$ 69,000 784 Yes Image
-GLOVE $\mathcal{X}_{GLOVE}$ 1,000,000 100 No Text
-NYtimes $\mathcal{X}_{NYtimes}$ 290,000 256 No Text
-Table 1: Descriptive traits characterising the datasets included in experiments.
+The reduced version of the NYtimes employed on this work was generated by building tf-idf descriptors from the bag-of-words version and embedding them into a lower dimensional space using the Johnson-Lindenstrauss Transform [67].
 4 Experimental results
-In order to show the versatility of GMASK, we present the analysis of four different datasets: municipalities, MNIST, GLOVE and NYtimes.
-This study consists of a neighbours search applying Euclidean distance, Manhattan distance and Chebyshev distance.
-The main objective is to measure the accuracy of the GMASK algorithm with different distances and to observe if any distance is more adequate for some dataset, depending on its characteristics.
-For each dataset, we compare the results obtained by GMASK with two other algorithms: PYNN and FLANN.
-Both PYNN and GMASK are able to work with these three distances, however, FLANN does not work with Chebyshev distance.
-Since MASK is a specific case of GMASK only able to work with Euclidean distance, its inclusion in these experiments has not been considered.
-Our approach to evaluating GMASK performance is conceived to closely emulate real-world conditions.
-The datasets used in this work are not artificially generated, and no preprocessing has been applied.
-The transformation and cleaning of the collected data is limited to dividing the original dataset into two distinct subgroups, the training and testing datasets.
-We also check that identical elements do not appear in both subgroups, thereby intensifying the complexity in the search for $k$ approximate neighbours.
+In order to prove the versatility of this framework, we have designed a series of experiments that may be repeated for every dataset described in Section 3.6.
+All of them have been randomly pre-split into train/test sets which, thereby intensifying the complexity, will not contain any common elements.
+First, by using the train set, we approach MSA using Euclidean, Manhattan, Chebyshev, and cosine distances, respectively.
+Then, by using NSA we perform searches for the 5, 10, and 15 approximate nearest neighbours of every query point contained in the corresponding test set.
+To evaluate the performance of GDASC, we compute the average recall of the searches conducted using each distance metric and compare the results with those obtained by employing two other algorithms: PyNN and FLANN.
+Our framework enables the use of all the distances proposed for these experiments, similar to PyNN.
+However, FLANN does not support Chebyshev or cosine distances.
+Once the average recall for each dataset-distance-method combination is obtained, we can determine if any specific distance metric is more suitable for a particular dataset based on its characteristics.
+This analysis supports one of the main hypotheses of this research: data indexing using non-Euclidean distances is not only feasible but also has the potential to yield improved outcomes.
 All experiments have been performed on a server equipped with 2 AMD EPYC 7451 microprocessors (24 cores/48 threads, 2.30 GHz, 64 MB L3 cache), 128 GB of DDR4 RAM and an Intel D3-S4510 SSD (capacity 480 GB) for secondary storage.
 The experimental results are presented in the following sections.
+(a) $\mathcal{X}_{muni}$ dataset
+(b) $\mathcal{X}_{MNIST}$ dataset
+(c) $\mathcal{X}_{GLOVE}$ dataset
+(d) $\mathcal{X}_{NYtimes}$ dataset
+Figure 6: Point plot illustrating the Average Recall of three different algorithms across various distances (manhattan, euclidean, chebyshev, cosine) in approximate nearest neighbour search for each dataset
 4.1 Municipalities
-The municipalities dataset is an interesting case due to the geographical distribution of points.
-In Figure 3, we can identify three main groups of points: the largest region corresponds to the peninsular territory of Spain, whereas the group of points located at the right of the map corresponds to the Balearic Islands and the group of points located at the bottom left corner of the map corresponds to the Canary Islands.
-The last two groups of points increase the difficulty of the analysis because they are far from the rest of the points, so we can consider them as outliers.
-Despite this dataset containing additional information, we only consider two variables: latitude and longitude.
-Figure 4 shows the recall comparison between FLANN, PYNN and GMASK.
-We use a heatmap to represent the recall value obtained for each algorithm with each distance in the search of 5, 10 and 15 neighbours.
-The recall obtained in this case for each method is very similar, and all of them depict good recall values, over 84%.
-It is noticeable that FLANN recall decreases by about 10% with the number of neighbours.
-Regardless of the distance used and the number of neighbours intended to search, GMASK is the only method that provides a 100% recall in all tests.
-Figure 4: Heatmap comparing the Recall (percentage of true nearest neighbours found) obtained by each algorithm and distance in approximate nearest neighbour search, for different number of neighbours, with dataset $\mathcal{X}_{muni}$.
+Figure 6a corresponds to a point plot that shows the average recall obtained in the $k$ approximate neighbours search with the FLANN, PyNN and GDASC methods, using Manhattan, Euclidean, Chebyshev and cosine distances.
+Regarding the GDASC algorithm, the structure generated by the MSA module has been built by setting the group length ($gl$) to 60 points, with 30 prototype points calculated for each group at the lowest level ($n_{1}$).
+Meanwhile, the NSA module uses a radius ($r$) of 30 to find the $k$ nearest neighbours within the structure created by the MSA module.
+For PYNNDescent and FLANN, the default parameter settings have been applied.
+For this dataset, the average recall achieved using Minkowski distances is considerably high for all methods.
+In particular, PyNN and GDASC achieve almost 100% recall in these cases.
+However, while GDASC maintains near-perfect performance with the cosine distance, the average recall of PyNN drops dramatically to 54%, yielding the worst result for this dataset.
 4.2 MNIST
-Given its high dimensionality and sparsity, MNIST has become a widespread benchmark in similarity search research.
-The following experiment evaluates the recall obtained in searching neighbours in this dataset, applying Manhattan, Euclidean and Chebyshev distance with FLANN, PYNN and GMASK methods.
-Figure 5 presents the results of this analysis.
-From this comparison, we can deduce that the results are highly comparable in the recall when applying PYNN and GMASK methods with Euclidean and Manhattan distances, respectively.
-Therefore, the combination of both distances and these methods suits this case study.
-However, this is not the case for Chebyshev distance since the recall is lower than 70%.
-There is also a meaningful reduction in FLANN recall when the number of neighbours increases.
-Regardless of the distance, GMASK yields a higher average recall than other methods.
-Figure 5: Heatmap comparing the Recall (percentage of true nearest neighbours found) obtained by each algorithm and distance in approximate nearest neighbour search, for different number of neighbours, with dataset $\mathcal{X}_{MNIST}$.
+MNIST is an interesting case of study due to its high dimensionality and sparsity.
+The results obtained in this experiment are presented in Figure 6b.
+Regarding the GDASC algorithm, the MSA module is initialised with parameters $gl=1000$ and $n_{1}$$prototypes=500$, while NSA has a $r=80000$ settled.
+For PYNNDescent and FLANN, default parameter settings have been applied.
+The analysis reveals that the average recall achieved by PyNN and GDASC are highly comparable, with GDASC consistently achieving slightly higher values.
+Both methods reach high recall values for all distances except for Chebyshev, which has an average recall below 70%.
+Additionally, the usefulness of FLANN is questioned not only by its lack of support for Chebyshev and cosine distances, but also by its low recall (less than 60%) for this dataset, which is more complex than the one analysed previously.
 4.3 GLOVE
-Hereafter, the third experiment involves the analysis of the GLOVE dataset, a high dimensional datasets that contains the largest number of elements among the datasets included in these experiments.
-Figure 6 depicts the recall results obtained for each method and distance.
-Once again, the luck of efficiency is evident in FLANN´s results.
-The best recall obtained with FLANN is 12% when searching for 5 neighbours applying Euclidean distance.
-PYNN method provides good accuracy when Manhattan distance is used.
-If we focus on PYNN method and compare the recall for the three distances, we can see that Manhattan distance provides the best results and Chebyshev distance provides the worst recalls.
-However, in the case of GMASK, the recall is 100% per each distance and with independence of the number of neighbours searched.
-Figure 6: Heatmap comparing the Recall (percentage of true nearest neighbours found) obtained by each algorithm and distance in approximate nearest neighbour search, for different number of neighbours, with dataset $\mathcal{X}_{GLOVE}$.
+Hereafter, the third experiment involves the analysis of the GLOVE dataset, a high-dimensional dataset that contains the largest number of elements among the datasets included in these experiments.
+Regarding GDASC, the MSA module is initialised with parameters $gl=1000$ and $n_{1}$$prototypes=500$, while NSA is launched with $r=250$.
+As previously, default parameter settings have been applied for PYNNDescent and FLANN.
+Figure 6c depicts the average recall performed by each method and the distance evaluated.
+Once again, the lack of efficiency is evident in FLANN´s results, which is not able to reach 15% with either the Euclidean or the Manhattan distance.
+If we focus on PyNN method and compare the recall for the four distances, we can see that the Manhattan distance provides a good accuracy and the Chebyshev distance gives the worst recall, while Euclidean and cosine distances provide a similar performance of over 60%.
+However, in the case of GDASC, it maintains a consistent average recall of 100%, regardless of the distance metric used.
 4.4 NYtimes
 Finally, the last experiment replicates the same analysis using a dataset characterised by high dimensionality, featuring a data-to-dimensionality ratio lower than that of GLOVE but higher than MNIST: the NYtimes dataset.
-Figure 7 illustrates the recall accomplished by FLANN, PYNN and GMASK.
-There is a noticeable difference between GMASK and the other methods, since the recall values for FLANN and PYNN are below 50%.
-However, GMASK consistently outperforms the other methods regardless of the distance chosen.
+In this scenario, GDASC MSA module is initialised with parameters $gl=1000$ and $n_{1}$$prototypes=500$ once again, while NSA is launched with $r=30$.
+For PYNNDescent and FLANN, default parameter settings have been applied.
+Figure 6d illustrates the average recall achieved by FLANN, PyNN and GDASC, highlighting the most striking finding of this experiment: achieving good performance depends not only on the use of a suitable method, but also on the choice of an appropriate distance metric.
+Although GDASC consistently outperforms the other two algorithms in all scenarios, the trends observed for PyNN are similar to GDASC.
+Both algorithms exhibit the lowest recall with the Euclidean distance, followed by the Manhattan distance, and show considerable improvement with the Chebyshev distance.
+The best results are obtained with the cosine distance, where GDASC achieves an average recall close to 100%, significantly overtaking PyNN, which reaches an average recall slightly above 65%.
 Once again, FLANN exhibits the poorest performance.
-Therefore, while the most remarkable insight obtained from this experiment is the superior performance of the Chebyshev distance metric compared to others, it is crucial to highlight that, in this case, GMASK nearly doubles the recall provided by PYNN, achieving an 82% recall when searching for the 5 nearest neighbours.
-Through this experiment, we empirically demonstrate what was previously stated theoretically in earlier sections of this manuscript: adopting the Euclidean distance as the default option in indexing structures and algorithms is a self-imposed restriction that can lead to sub-optimal results.
-Figure 7: Heatmap comparing the Recall (percentage of true nearest neighbours found) obtained by each algorithm and distance in approximate nearest neighbour search, for different number of neighbours, with dataset $\mathcal{X}_{NYtimes}$.
+Through this experiment, we empirically demonstrate what has been stated theoretically in earlier sections of this manuscript: adopting Euclidean distance as the default option in indexing structures and algorithms is a self-imposed restriction that can lead to suboptimal results.
 5 Discussion
-Given the notable prevalence of Euclidean distance in similarity searches in the previous literature, as well as the limitations detected in the widely adopted $k$-means algorithm, which is strongly linked to this metric, there is a growing need to explore the use of alternative metrics when addressing these types of problems.
-Stemming from this need, GMASK makes the accommodation of arbitrary distances feasible by integrating the $k$-medoids algorithm as a clustering algorithm, allowing us to leverage the advantages derived from using several distance metrics on high-dimensional datasets.
-It is crucial to emphasise that the algorithm’s generalisation, achieved through the integration of $k$-medoids, significantly broadens the range of applicable distances while preserving the structural integrity and properties of the original algorithm.
-Thus, GMASK maintains crucial characteristics, including the desirable scalability property facilitated by the bottom-top construction of the search tree, retaining a competitive edge over other state-of-the-art algorithms that require centralising all data on a single server.
-As elaborated in Section “Experimental results”, our findings demonstrate the efficiency of this new approach, thereby supporting the hypothesis that alternative distances beyond the Euclidean metric may yield superior results in solving nearest-neighbour queries, contingent upon the specific characteristics of the dataset.
-Therefore, it becomes evident that the outcomes achieved using the proposed method surpass those obtained through two other widely adopted state-of-the-art algorithms, which are distinct from each other and significantly differ from GMASK.
+Given the notable prevalence of Euclidean distance in similarity searches in the previous literature, as well as the limitations detected in the widely adopted $k$-means algorithm, strongly linked to this metric, there is a growing need to explore the use of alternative metrics when addressing these types of problems[51].
+Therefore, this paper introduces GDASC, a versatile data indexing framework that facilitates the incorporation of arbitrary distances by employing different clustering algorithms that assign a representative element to each cluster and allow the use of multiple $\ell_{p}$ norms or even other distance functions.
+This approach significantly broadens the spectrum of compatible problems while offering desirable scalability through the bottom-up construction of the search tree, providing a competitive edge over other state-of-the-art algorithms that require centralisation of all data on a single server.
+As detailed in Section 4, experimental results using $k$-medoids clustering in GDASC with real datasets varying in type, size, dimensionality, and sparsity confirm the applicability of this approach for approximate similarity search, with outcomes surpassing those obtained with two other widely adopted MAM that differ significantly from GDASC.
+Our findings not only demonstrate the efficiency of this novel framework but also support the hypothesis that the use of alternative distances beyond the Euclidean metric may yield superior results for nearest-neighbour queries, depending on the specific characteristics of the dataset.
 These findings would also be in line with previous publications [46] suggesting that, for a given high-dimensional problem, using an $\ell_{p}$ norm with lower values of $p$ may be advantageous in assessing dissimilarity between query points.
@@
 However, further empirical research is required to validate this observed behaviour more precisely.
-GMASK provides an avenue for future research to integrate distance functions beyond Minkowski, thus providing many opportunities to extend this research line.
-For instance, it would be interesting to explore the suitability of distance functions for specific purposes, such as cosine similarity for textual data or the Jaccard distance for categorical data, to address particular problems.
-Furthermore, although GMASK has yielded notably satisfactory experimental results, there still exists potential for enhancing this implementation to achieve further optimisation in computational methods, thereby increasing overall performance.
-To this end, the proposal suggests exploring the incorporation of a just-in-time (JIT) compiler for Python, such as Numba  [^8] https://numba.pydata.org—employed by other algorithms such as PYNNDescent—, or considering the adaptation and port of GMASK to a language renowned for its strong performance in building scalable and concurrent systems like Go  [^9] https://go.dev.
+GDASC provides an avenue for future research to integrate distance functions beyond Minkowski or cosine similarity, thus providing many opportunities to extend this research line.
+For instance, it would be interesting to explore the suitability of distance functions for specific purposes, such as the Jaccard distance for categorical data, to address particular problems.
+Furthermore, although GDASC has yielded notably satisfactory experimental results, there still exists potential for enhancing this implementation to achieve further optimisation in computational methods, thereby increasing overall performance.
+To this end, the proposal suggests exploring the incorporation of a just-in-time (JIT) compiler for Python, such as Numba  [^8] https://numba.pydata.org—employed by other algorithms such as PyNN—, or considering the adaptation and port of GDASC to a language renowned for its strong performance in building scalable and concurrent systems like Go  [^9] https://go.dev.
 6 Conclusions and future work
@@
 In many cases, these datasets must be stored in distributed systems to become manageable, imposing additional restrictions that rule out algorithms conceived for cases where data must reside on a single machine.
-This paper presents the GMASK algorithm for approximate similarity search in distributed systems that can accept arbitrary distance functions.
-One of the main novelties is adopting the $k$-medoids (PAM) algorithm, which is not bounded to the Euclidean distance, to achieve efficient solutions.
-Furthermore, using FastPAM, a highly efficient implementation of the PAM algorithm to solve the $k$-medoids problem, GMASK can improve its efficiency with respect to other previous algorithms.
-Experiments conducted on real-world datasets with different properties confirm that GMASK, based on $k$-medoids, performs better in solving nearest-neighbour queries than other frequently used algorithms, using other Minkowski dissimilarity functions different from the Euclidean distance.
+This paper presents the GDASC algorithm for approximate similarity search in distributed systems that can accept arbitrary distance functions.
+One of the main novelties is the integration of $k$-medoids as clustering method, which is not bounded to the Euclidean distance, to achieve efficient solutions.
+Experiments conducted on real-world datasets with different properties confirm that GDASC performs better in solving nearest-neighbour queries than other frequently used algorithms, using dissimilarity functions different from the Euclidean distance.
 These results illustrate that indexing data with non-Euclidean distances is not only viable but also holds the potential to enhance search outcomes substantially.
 Overlooking alternative distances would represent a voluntary limitation that may hinder us from achieving the optimal results.
-GMASK lays out new interesting routes for further research.
-Once the requirement of using the Euclidean distance has been lifted, additional studies should be conducted to evaluate GMASK performance in other settings where alternative distances can be applied, such as pseudo-metrics (Minkowski distances with $p<1$) or non-metric dissimilarities.
+GDASC lays out new interesting routes for further research.
+Once the requirement of using the Euclidean distance has been lifted, additional studies should be conducted to evaluate GDASC performance in other settings where alternative distances can be applied, such as pseudo-metrics (Minkowski distances with $p<1$) or non-metric dissimilarities.
 Besides, using arbitrary distances will enable us to consider datasets with qualitative features exclusively or a combination of quantitative and qualitative features.
 Many state-of-the-art algorithms and standard dissimilarity measures cannot be directly applied in these cases since they are incompatible with categorical features.
+7 Availability of data and material
+7.1 Data availability
+All datasets used in experiments are publicly available:
+7.2 Code availability
+All code related to the proposed algorithm and the implementation of experiments is hosted in the elenagarciamorato/GDASC GitHub repository, specifically in the release with DOI 10.5281/zenodo.13220360).